### 02 Apr 2026 v1

# A Digit-Sum Formula for Böttcher Coordinates

Abstract.
Let $p$ be an odd prime and let $\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}$. We prove an unconditional digit-sum congruence formula for the coefficients of the Böttcher coordinate of $\phi$. The argument separates the divisible case $k\equiv 0\pmod{p}$ from the general case, gives a complete vector-partition description of the surviving $B$-term monomials, and reduces the remaining partition sum to a one-variable cumulant. As consequences we recover the residue-class congruences conjectured by Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$.
Key words and phrases:
Böttcher coordinate, superattracting germ, non-archimedean dynamics
2020 Mathematics Subject Classification:
37P35, 37F50, 11S82
1. Introduction
The local study of holomorphic iteration begins with the functional equations of Schröder and Koenigs, and the theory of complex dynamics took its modern shape through the early twentieth-century work of Böttcher, Fatou, and Julia [1, 6]. For a one-dimensional analytic germ with a superattracting fixed point, Böttcher’s theorem provides a canonical local coordinate in which the map becomes a pure monomial. In polynomial dynamics this coordinate at infinity does considerably more than linearize the map: it identifies the basin of infinity with the exterior disk, encodes the escape-rate function, and underlies external rays, landing theory, and the combinatorial study of Julia sets and the Mandelbrot set [2, 3, 6]. The same normal form also appears in several variables, where the local geometry near superattracting fixed points becomes subtler [5].
In recent years Böttcher coordinates have also become important in non-archimedean dynamics. Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs over $p$-adic rings [8], while Fu and Nie later settled a conjecture of Salerno–Silverman on the radius of convergence in the wild superattracting case [4]. The family relevant for the present paper is
$\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}.$
If
$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k}$
denotes the Böttcher coordinate characterized by
$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{p^{2}}),$
then Salerno and Silverman recorded striking mod-$p$ patterns in the coefficients $a_{k}$; their original formulation is Conjecture 25 of [8]. For an integer $n=\sum_{i\geq 0}n_{i}p^{i}$ with $0\leq n_{i}\leq p-1$, let
$S_{p}(n):=\sum_{i\geq 0}n_{i}$
be the sum of its base-$p$ digits.
From now on, we assume that $p\geq 3$. We note that the corresponding result for $p=2$ can be established similarly, but since the arguments for all $p\geq 3$ fit into a unified framework, we do not treat the case $p=2$ in the present paper. The following are our main theorems.
Theorem 1.1 (Digit-sum formula).
Let $k\geq 1$ and write
$k=a+d_{1}p+d_{2}p^{2}+\cdots+d_{r}p^{r},\qquad 0\leq a,d_{i}\leq p-1,$
with
$s:=d_{1}+\cdots+d_{r}=S_{p}(k)-a.$
Then
(1.1) $a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
In fact, we have a stronger result
(1.2) $a_{k}\equiv(-1)^{S_{p}(k)}(a+1)^{S_{p}(k)-1}\pmod{p}$.
We also recover the original residue-class conjecture of Salerno and Silverman.
Theorem 1.2 (Salerno–Silverman’s Conjecture 25).
For every integer $m\geq 1$ one has
$a_{pm}\equiv(-1)^{m},\qquad a_{pm-1}\equiv 0,\qquad a_{pm-2}\equiv-1\pmod{p}$.
The proof is organized by separating the residue class $a=0$ from the classes $a\neq 0$. Section 2 recalls the AABBCC recursion and proves the basic $A$- and $C$-term lemmas. Section 3 supplies a complete proof of the divisible case $a=0$. The remaining sections treat $a\neq 0$: Section 4 gives a vector-partition expansion of the $B$-term, Section 5 evaluates the first block of coefficients and collapses the multivariate cumulant, and Section 6 proves (1.1). In the final section we prove (1.2) and deduce Theorem 1.2.
2. The recursion and two structural lemmas
Let $q:=p^{2}$. As in [8, §5], write
$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k},\qquad a_{0}=1,$
so that
$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{q}).$
The coefficient comparison may be organized in the form
(2.1) $a_{k}=A_{k}[x^{k}]-B_{k}[x^{k}]-C_{k}[x^{k}],$
where
$A_{k}:=\frac{k!}{q}\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{q\ell},\qquad B_{k}:=\frac{k!}{q}\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q},\qquad C_{k}:=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q+1}.$
Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. Throughout we use Legendre’s formula
(2.2) $\operatorname{ord}_{p}(n!)=\frac{n-S_{p}(n)}{p-1}$
and the standard digit-sum subadditivity
(2.3) $S_{p}(u+v)\leq S_{p}(u)+S_{p}(v)$
for nonnegative integers $u,v$, together with its obvious extension to finite sums.
Proposition 2.1 (The $A$-term always vanishes modulo $p$).
For every $k\geq 1$ one has
$A_{k}[x^{k}]\equiv 0\pmod{p}$.
Proof.
If $q\nmid k$, then $A_{k}[x^{k}]=0$ because every monomial occurring in $A_{k}$ has degree divisible by $q=p^{2}$.
Assume now that $k=qn=p^{2}n$. Then only the term $\ell=n$ contributes, so
$A_{k}[x^{k}]=\frac{(qn)!}{q\,n!}\,a_{n}=\frac{(qn)!}{p^{2}n!}\,a_{n}.$
Since multiplying by $p^{2}$ only shifts the base-$p$ digits, we have $S_{p}(qn)=S_{p}(n)$. Hence
$\operatorname{ord}_{p}\!\left(\frac{(qn)!}{p^{2}n!}\right)$ $=\operatorname{ord}_{p}((qn)!)-2-\operatorname{ord}_{p}(n!)$
$=\frac{qn-S_{p}(qn)}{p-1}-2-\frac{n-S_{p}(n)}{p-1}$
$=\frac{(p^{2}-1)n}{p-1}-2$
$=(p+1)n-2.$
Since $n\geq 1$ and $p\geq 3$, this is at least $1$. Therefore $p\mid A_{k}[x^{k}]$, so $A_{k}[x^{k}]\equiv 0\pmod{p}$. ∎
For the $B$-term, let $a^{e}=\prod_{n\geq 0}a_{n}^{e_{n}}$ be a monomial occurring in $B_{k}[x^{k}]$. Then
$\sum_{n\geq 0}e_{n}=q,\qquad\sum_{n\geq 0}ne_{n}=k,$
and the coefficient of $a^{e}$ in $B_{k}[x^{k}]$ is
$B_{k}[x^{k}][a^{e}]=\frac{k!}{q}\binom{q}{e_{0},e_{1},\dots}\prod_{n\geq 0}\frac{1}{(n!)^{e_{n}}}.$
A direct application of Legendre’s formula yields
(2.4) $(p-1)\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=-2p+1-S_{p}(k)+\sum_{n\geq 0}S_{p}(e_{n})+\sum_{n\geq 0}e_{n}S_{p}(n).$
Equivalently, with
$c(e):=\frac{\sum_{n\geq 0}e_{n}S_{p}(n)-S_{p}(k)}{p-1},\qquad d(e):=\frac{\sum_{n\geq 0}S_{p}(e_{n})-1}{p-1},$
one has
(2.5) $\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=c(e)+d(e)-2.$
The two quantities $c(e)$ and $d(e)$ measure the carry defects in the base-$p$ additions $\sum_{n}e_{n}n=k$ and $\sum_{n}e_{n}=p^{2}$, respectively.
Proposition 2.2 (A global $C$-term lemma).
If $k\equiv a\pmod{p}$ with $1\leq a\leq p-1$, then
$C_{k}[x^{k}]\equiv a\,a_{k-1}\pmod{p}$.
If $p\mid k$, then
$C_{k}[x^{k}]\equiv 0\pmod{p}$.
Proof.
Let $a^{e}=\prod_{n=0}^{k-1}a_{n}^{e_{n}}$ be a monomial occurring in $C_{k}[x^{k}]$. Since
$C_{k}=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{p^{2}+1},$
its monomials satisfy
$\sum_{n=0}^{k-1}e_{n}=p^{2}+1,\qquad\sum_{n=0}^{k-1}ne_{n}=k-1.$
The coefficient of $a^{e}$ in $C_{k}[x^{k}]$ is
$C_{k}[x^{k}][a^{e}]=k!\binom{p^{2}+1}{e_{0},e_{1},\dots,e_{k-1}}\prod_{n=0}^{k-1}\frac{1}{(n!)^{e_{n}}},$
and a direct Legendre computation gives
(2.6) $(p-1)\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)=-S_{p}(k)-1+\Sigma(e)+T(e),$
where
$\Sigma(e):=\sum_{n=0}^{k-1}S_{p}(e_{n}),\qquad T(e):=\sum_{n=0}^{k-1}e_{n}S_{p}(n).$
Assume first that $p\mid k$. Write $k=p^{\nu}u$ with $\nu\geq 1$ and $p\nmid u$. Subtracting $1$ from $k$ changes the final $\nu$ zeros in the base-$p$ expansion into p1p-1’s and lowers the next digit by $1$, so
$S_{p}(k-1)=S_{p}(k)-1+\nu(p-1).$
Also,
$\Sigma(e)\geq S_{p}(p^{2}+1)=2,\qquad T(e)\geq S_{p}(k-1).$
Therefore every monomial coefficient in $C_{k}[x^{k}]$ satisfies
$(p-1)\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)$ $\geq-S_{p}(k)-1+2+S_{p}(k-1)$
$=\nu(p-1)\geq p-1.$
Hence $\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)\geq 1$ for every monomial, and so $C_{k}[x^{k}]\equiv 0\pmod{p}$.
Now assume that $k\equiv a\pmod{p}$ with $1\leq a\leq p-1$. Then
$S_{p}(k-1)=S_{p}(k)-1.$
Since again $\Sigma(e)\geq 2$ and $T(e)\geq S_{p}(k-1)$, every monomial that survives modulo $p$ must satisfy
(2.7) $\Sigma(e)=2,\qquad T(e)=S_{p}(k-1).$
The condition $\Sigma(e)=2$ leaves exactly two multiplicity patterns:
$e_{t}=p^{2}+1\qquad\text{or}\qquad e_{r}=p^{2},\ e_{s}=1,$
with all other multiplicities equal to 0.
In the first case,
$k-1=(p^{2}+1)t,\qquad T(e)=(p^{2}+1)S_{p}(t).$
If $t>0$, then
$S_{p}(k-1)=S_{p}((p^{2}+1)t)\leq S_{p}(p^{2}t)+S_{p}(t)=2S_{p}(t)<(p^{2}+1)S_{p}(t)=T(e),$
contradicting (2.7). Hence $t=0$, so $k=1$, and the claimed congruence is immediate.
In the second case,
$k-1=p^{2}r+s,\qquad T(e)=p^{2}S_{p}(r)+S_{p}(s).$
If $r>0$, then
$S_{p}(k-1)=S_{p}(p^{2}r+s)\leq S_{p}(r)+S_{p}(s)<p^{2}S_{p}(r)+S_{p}(s)=T(e),$
again a contradiction. Therefore $r=0$, so s=k1s=k-1 and the unique surviving monomial is
$e_{0}=p^{2},\qquad e_{k-1}=1.$
Its coefficient is
$k!\binom{p^{2}+1}{p^{2},1}\frac{1}{(k-1)!}=k(p^{2}+1)\equiv k\equiv a\pmod{p}$.
Thus
$C_{k}[x^{k}]\equiv a\,a_{k-1}\pmod{p}$,
as claimed. ∎
Theorem 2.3 (The case $a=0$).
For every integer $M\geq 1$ one has
$a_{pM}\equiv(-1)^{M}\pmod{p}$.
The proof of Theorem 2.3 is given in the next section.
3. The Case $a=0$
This section gives a complete proof of Theorem 2.3. The main new input is an explicit coefficient identity for a truncated exponential series.
Lemma 3.1 (A truncated-exponential coefficient identity).
For $k\geq 1$, define
$S_{k-1}(y):=\sum_{\ell=0}^{k-1}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
Then
(3.1) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=[y^{k}]e^{-qy/p}-q\frac{(-1)^{k}}{p^{k}k!}.$
Equivalently,
(3.2) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=\frac{(-q/p)^{k}}{k!}-q\frac{(-1)^{k}}{p^{k}k!}.$
Since $q=p^{2}$, this becomes
(3.3) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=\frac{(-p)^{k}}{k!}-q\frac{(-1)^{k}}{p^{k}k!}.$
Proof.
From the exponential series,
$e^{-y/p}=\sum_{\ell=0}^{\infty}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
Hence
$S_{k-1}(y)=e^{-y/p}-R_{k}(y),\qquad R_{k}(y):=\sum_{\ell=k}^{\infty}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
In particular,
$R_{k}(y)=\frac{(-1)^{k}}{p^{k}k!}y^{k}+O(y^{k+1}),$
so
$S_{k-1}(y)=E(y)+U(y),\qquad E(y):=e^{-y/p},\qquad U(y):=-\frac{(-1)^{k}}{p^{k}k!}y^{k}+O(y^{k+1}).$
Now expand the $q$th power:
$(E+U)^{q}=E^{q}+qE^{q-1}U+\sum_{j=2}^{q}\binom{q}{j}E^{q-j}U^{j}.$
Since $\operatorname{ord}_{y}(U)\geq k$, every term with $j\geq 2$ has $y$-adic order at least $2k>k$, hence contributes nothing to the coefficient of $y^{k}$. Therefore
$y^{k}](E+U)^{q}=[y^{k}]E^{q}+[y^{k}](qE^{q-1}U).$
Because $E(0)=1$, we have $E^{q-1}(y)=1+O(y)$, and hence
$y^{k}](E^{q-1}U)=[y^{k}]U=-\frac{(-1)^{k}}{p^{k}k!}.$
Substituting gives
$y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=[y^{k}]E(y)^{q}-q\frac{(-1)^{k}}{p^{k}k!}.$
Finally, $E(y)^{q}=e^{-qy/p}$, which proves (3.1); equation (3.2) follows immediately from the exponential series. ∎
Corollary 3.2.
For every $m\geq 1$,
$y^{m}]\left(\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}\right)^{q}=\frac{p^{m}}{m!}-\frac{q}{p^{m}m!}.$
Proof.
Apply Lemma 3.1 with $k=m$ and replace $y$ by y-y. Since
$\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}=S_{m-1}(-y),$
we obtain
$y^{m}]\left(\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}\right)^{q}=[y^{m}]e^{qy/p}-q\frac{1}{p^{m}m!}.$
Because $q=p^{2}$, we have $e^{qy/p}=e^{py}$, whose $y^{m}$-coefficient is $p^{m}/m!$. ∎
Proposition 3.3 (Structure of the surviving $B$-term monomials).
Assume that $p\mid k$. Write
$B_{k}[x^{k}]=\sum_{e}\gamma_{B}(e)a^{e},$
where the sum is over all multi-indices $e=(e_{0},\dots,e_{k-1})$ satisfying
$\sigma(e):=\sum_{n=0}^{k-1}e_{n}=q,\qquad\nu(e):=\sum_{n=1}^{k-1}ne_{n}=k,\qquad a^{e}:=\prod_{n=0}^{k-1}a_{n}^{e_{n}}.$
Suppose that a coefficient $\gamma_{B}(e)$ has $p$-adic valuation 0. Then either
$a^{e}=a_{0}^{q-p}a_{1}^{p},$
or else every positive index $n$ with $e_{n}>0$ is divisible by $p$.
Proof.
For such $e$, the coefficient of $a^{e}$ in $B_{k}[x^{k}]$ is
$\gamma_{B}(e):=\frac{k!}{q}\binom{q}{e_{0},\dots,e_{k-1}}\prod_{n=0}^{k-1}\frac{1}{(n!)^{e_{n}}}.$
A straightforward calculation with Legendre’s formula gives
(3.4) $(p-1)\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=-2p+1-S_{p}(k)+\sum_{n=0}^{k-1}\bigl(S_{p}(e_{n})+e_{n}S_{p}(n)\bigr).$
It is convenient to write
$c(e):=\frac{\sum_{n=1}^{k-1}e_{n}S_{p}(n)-S_{p}(k)}{p-1},\qquad d(e):=\frac{\sum_{n=0}^{k-1}S_{p}(e_{n})-1}{p-1}.$
By digit-sum subadditivity, both numerators are nonnegative; since $S_{p}(N)\equiv N\pmod{p-1}$ for every integer $N$, both numerators are divisible by p1p-1, so $c(e),d(e)\in\mathbf{Z}_{\geq 0}$. Using
$\sum_{n=0}^{k-1}e_{n}=q=p^{2},\qquad S_{p}(q)=1,$
we can rewrite (3.4) as
(3.5) $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)+d(e)-2.$
Now assume that $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=0$.
Step 1: if some occurring positive index is not divisible by $p$, then $c(e)\geq 1$. Suppose there exists $n\geq 1$ with $e_{n}>0$ and $p\nmid n$. In the base-$p$ sum
$\sum_{n=1}^{k-1}e_{n}n=k,$
the units digit on the left receives a nonzero contribution from such an $n$, while the right-hand side has units digit 0 because $p\mid k$. Therefore at least one carry must occur from the units place to the $p$-place. Equivalently, the digit-sum defect
$\sum_{n=1}^{k-1}e_{n}S_{p}(n)-S_{p}(k)$
is positive, so $c(e)\geq 1$.
Step 2: $d(e)\neq 0$. Assume for contradiction that $d(e)=0$. Then
$\sum_{n=0}^{k-1}S_{p}(e_{n})=1.$
Since every $e_{n}\geq 0$ and $\sum e_{n}=q=p^{2}$, the only possibility is that one index $m$ satisfies $e_{m}=p^{2}$ and all other $e_{n}$ vanish. In that case,
$k=\sum_{n=1}^{k-1}ne_{n}=mp^{2},$
and therefore
$c(e)=\frac{p^{2}S_{p}(m)-S_{p}(p^{2}m)}{p-1}=\frac{(p^{2}-1)S_{p}(m)}{p-1}=(p+1)S_{p}(m)\geq p+1.$
Hence (3.5) gives
$\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)-2\geq p-1>0,$
contradicting the assumption that the valuation is 0. Thus $d(e)\geq 1$.
Step 3: from $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=0$ we obtain $c(e)=d(e)=1$. By Step 1, whenever some occurring positive index is not divisible by $p$ we have $c(e)\geq 1$, and by Step 2 we always have $d(e)\geq 1$. Since (3.5) says
$0=\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)+d(e)-2,$
we must have
$c(e)=d(e)=1.$
Step 4: $d(e)=1$ forces every $e_{n}$ to be divisible by $p$. The condition $d(e)=1$ means that
$\sum_{n=0}^{k-1}S_{p}(e_{n})-S_{p}\!\left(\sum_{n=0}^{k-1}e_{n}\right)=p-1,$
so the addition of the integers $e_{0},e_{1},\dots,e_{k-1}$ to obtain $q=p^{2}$ produces exactly one carry in base $p$. Now suppose some $e_{n}$ has nonzero units digit. Then in the sum
$\sum_{n=0}^{k-1}e_{n}=p^{2},$
there is a carry from the units place to the $p$-place. But the final result $p^{2}$ has base-$p$ digits $(0,0,1)$, so in order for the $p$-place digit also to end up equal to 0, there must be a second carry from the $p$-place to the $p^{2}$-place. This contradicts the fact that there is only one carry. Hence every $e_{n}$ is divisible by $p$.
So we may write
$e_{n}=pf_{n},\qquad\sum_{n=0}^{k-1}f_{n}=p.$
Then
$k=\sum_{n=1}^{k-1}ne_{n}=p\sum_{n=1}^{k-1}f_{n}n,$
so
$S_{p}(k)=S_{p}\!\left(\sum_{n=1}^{k-1}f_{n}n\right)\leq\sum_{n=1}^{k-1}f_{n}S_{p}(n).$
Therefore
$c(e)=\frac{p\sum_{n=1}^{k-1}f_{n}S_{p}(n)-S_{p}(k)}{p-1}\geq\sum_{n=1}^{k-1}f_{n}S_{p}(n).$
Since $c(e)=1$, it follows that
$\sum_{n=1}^{k-1}f_{n}S_{p}(n)\leq 1.$
The left-hand side is a sum of nonnegative integers, and it is positive because $k>0$; thus exactly one positive index $n$ satisfies $f_{n}=1$, and for that index we have $S_{p}(n)=1$. Therefore $n$ is a power of $p$, and the positive part of the monomial is $a_{n}^{p}$.
If this $n$ is not divisible by $p$, then the only possibility is $n=1$, and since $\sum f_{n}=p$ we must have $f_{0}=p-1$, that is,
$e_{0}=p(p-1)=q-p,\qquad e_{1}=p.$
This gives the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$. In all other cases, every positive index that occurs is divisible by $p$. ∎
Corollary 3.4 (Dependence only on the earlier $a_{pi}$).
If $p\mid k$ and $k>p$, then modulo $p$ the quantity $a_{k}$ depends only on the earlier coefficients $a_{pi}$ with $pi<k$.
Proof.
By Propositions 2.1 and 2.2,
$a_{k}\equiv-B_{k}[x^{k}]\pmod{p}\qquad(p\mid k).$
Now Proposition 3.3 shows that any monomial in $B_{k}[x^{k}]$ whose coefficient can survive modulo $p$ is either the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$ or else involves only indices divisible by $p$. If $k>p$, then the exceptional monomial cannot occur, because its weight is
$0\cdot(q-p)+1\cdot p=p,$
not $k$. Therefore every surviving monomial in $B_{k}[x^{k}]$ involves only $a_{pi}$ with $pi<k$. ∎
Lemma 3.5.
The first nontrivial coefficient of $f_{\phi}$ is
$a_{1}=-1.$
Proof.
Write
$f_{\phi}(x)=x+a_{1}x^{2}+O(x^{3}).$
Then
$\phi(f_{\phi}(x))=f_{\phi}(x)^{q}+p^{2}f_{\phi}(x)^{q+1}.$
Expanding to order $x^{q+1}$ gives
$f_{\phi}(x)^{q}=x^{q}+qa_{1}x^{q+1}+O(x^{q+2}),\qquad p^{2}f_{\phi}(x)^{q+1}=p^{2}x^{q+1}+O(x^{q+2}).$
On the other hand,
$f_{\phi}(x^{q})=x^{q}+O(x^{2q}),$
so comparison of the $x^{q+1}$-coefficients in the Böttcher equation yields
$qa_{1}+p^{2}=0.$
Since $q=p^{2}$, we obtain $a_{1}=-1$. ∎
Lemma 3.6.
One has
$a_{p}\equiv-1\pmod{p}$.
Proof.
By Propositions 2.1 and 2.2,
$a_{p}\equiv-B_{p}[x^{p}]\pmod{p}$.
Since there are no positive indices less than $p$ that are divisible by $p$, Proposition 3.3 shows that the only monomial in $B_{p}[x^{p}]$ whose coefficient can be a $p$-adic unit is the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$. Its coefficient is
$\gamma_{p}=\frac{p!}{q}\binom{q}{q-p,p}=\frac{1}{p^{2}}\prod_{j=0}^{p-1}(p^{2}-j).$
Reducing modulo $p$ gives
$\gamma_{p}\equiv\prod_{j=1}^{p-1}(-j)=(-1)^{p-1}(p-1)!\equiv-1\pmod{p}$
by Wilson’s theorem. Using $a_{0}=1$ and Lemma 3.5, we find
$B_{p}[x^{p}]\equiv\gamma_{p}a_{1}^{p}\equiv(-1)(-1)^{p}\equiv 1\pmod{p}$.
Therefore
$a_{p}\equiv-1\pmod{p}$.
Proof of Theorem 2.3.
We argue by induction on $M$. The case $M=1$ is Lemma 3.6.
Assume now that $M\geq 2$ and that
(3.6) $a_{pi}\equiv(-1)^{i}\pmod{p}\qquad(1\leq i<M).$
Applying Corollary 3.4 with $k=pM$, we obtain
(3.7) $a_{pM}\equiv-B_{pM}[x^{pM}]\pmod{p}$,
and only the indices divisible by $p$ can contribute. Therefore, after setting $y=x^{p}$,
(3.8) $B_{pM}[x^{pM}]\equiv\frac{(pM)!}{q}[y^{M}]\left(\sum_{i=0}^{M-1}\frac{a_{pi}}{(pi)!}y^{i}\right)^{q}\pmod{p}$.
Here $a_{0}=1$ corresponds to $i=0$.
For $i\geq 0$, define
$\nu_{i}:=\frac{(pi)!}{p^{i}i!}.$
Then $\nu_{i}\in\mathbf{Z}$ and
(3.9) $\nu_{i}=\prod_{r=1}^{i}\prod_{t=1}^{p-1}\bigl(p(r-1)+t\bigr)\equiv((p-1)!)^{i}\equiv(-1)^{i}\pmod{p}$
by Wilson’s theorem. Hence, by the induction hypothesis (3.6),
$\omega_{i}:=a_{pi}\nu_{i}^{-1}\in\mathbf{Z}_{(p)}$
satisfies $\omega_{i}\equiv 1\pmod{p}$ for $0\leq i<M$ (with $\omega_{0}=1$). Consequently,
(3.10) $\frac{a_{pi}}{(pi)!}=\frac{\omega_{i}}{p^{i}i!}\qquad(0\leq i<M).$
Write $\omega_{i}=1+pz_{i}$ with $z_{i}\in\mathbf{Z}_{(p)}$, and define
$H_{M}(y):=\sum_{i=0}^{M-1}\frac{y^{i}}{p^{i}i!},\qquad L_{M}(y):=\sum_{i=0}^{M-1}z_{i}\frac{y^{i}}{p^{i}i!}.$
Then by (3.10),
$\sum_{i=0}^{M-1}\frac{a_{pi}}{(pi)!}y^{i}=H_{M}(y)+pL_{M}(y).$
We claim that
(3.11) $\frac{p^{M}M!}{q}[y^{M}]\bigl(H_{M}(y)+pL_{M}(y)\bigr)^{q}\equiv\frac{p^{M}M!}{q}[y^{M}]H_{M}(y)^{q}\pmod{p}$.
Indeed,
$\bigl(H_{M}+pL_{M}\bigr)^{q}-H_{M}^{q}=\sum_{j=1}^{q}\binom{q}{j}p^{j}H_{M}^{q-j}L_{M}^{j}.$
Any term contributing to the coefficient of $y^{M}$ in $H_{M}^{q-j}L_{M}^{j}$ has total denominator $p^{M}\prod i!^{e_{i}}$, so
$y^{M}]H_{M}^{q-j}L_{M}^{j}\in p^{-M}\mathbf{Z}_{(p)}.$
Therefore it is enough to show that
$\operatorname{ord}_{p}\!\left(\frac{p^{j}}{q}\binom{q}{j}\right)\geq 1\qquad(1\leq j\leq q).$
Using
$\binom{q}{j}=\frac{q}{j}\binom{q-1}{j-1}$,
we get
$\frac{p^{j}}{q}\binom{q}{j}=\frac{p^{j}}{j}\binom{q-1}{j-1}$,
so
$\operatorname{ord}_{p}\!\left(\frac{p^{j}}{q}\binom{q}{j}\right)\geq j-\operatorname{ord}_{p}(j)\geq 1.$
This proves (3.11).
Now combine (3.7), (3.8), and (3.11). Since $(pM)!=p^{M}M!\nu_{M}$, we obtain
(3.12) $a_{pM}\equiv-\nu_{M}\frac{p^{M}M!}{q}[y^{M}]H_{M}(y)^{q}\pmod{p}$.
By Corollary 3.2,
$y^{M}]H_{M}(y)^{q}=\frac{p^{M}}{M!}-\frac{q}{p^{M}M!}.$
Substituting into (3.12) gives
$a_{pM}\equiv-\nu_{M}\frac{p^{M}M!}{q}\left(\frac{p^{M}}{M!}-\frac{q}{p^{M}M!}\right)\equiv-\nu_{M}\bigl(p^{2M-2}-1\bigr)\pmod{p}$.
Since $M\geq 2$, we have $p^{2M-2}\equiv 0\pmod{p}$, so
$a_{pM}\equiv\nu_{M}\pmod{p}$.
Finally, (3.9) gives $\nu_{M}\equiv(-1)^{M}\pmod{p}$, and hence
$a_{pM}\equiv(-1)^{M}\pmod{p}$.
This completes the induction. ∎
4. Vector partitions and the $B$-term for $a\neq 0$
The higher base-$p$ digits of the index will be encoded by vectors.
Definition 4.1.
Fix $r\geq 1$. A digit vector is an element
$d=(d_{1},\dots,d_{r})\in\{0,1,\dots,p-1\}^{r}.$
Its weight is
$|d|:=d_{1}+\cdots+d_{r},$
and its numeric value is
$N(d):=\sum_{i=1}^{r}d_{i}p^{i}.$
A vector partition of $d$ is a formal product
$\Pi=\prod_{\beta\neq 0}\beta^{r_{\beta}},$
with $r_{\beta}\in\mathbf{Z}_{\geq 0}$ and
$\sum_{\beta\neq 0}r_{\beta}\beta=d.$
The total number of blocks is
$R(\Pi):=\sum_{\beta\neq 0}r_{\beta}.$
Fix $a\in\{0,1,\dots,p-1\}$. For a digit vector $\beta$ define
$H_{\beta}(y):=\sum_{j=0}^{a}\frac{a_{N(\beta)+j}}{j!}y^{j},\qquad F_{a}(y):=\sum_{j=0}^{a}\frac{a_{j}}{j!}y^{j},\qquad\widetilde{H}_{d}(y):=\sum_{j=0}^{a-1}\frac{a_{N(d)+j}}{j!}y^{j}.$
Write also
$d!:=\prod_{i=1}^{r}d_{i}!,\qquad\beta!:=\prod_{i=1}^{r}\beta_{i}!\qquad\text{for }\beta=(\beta_{1},\dots,\beta_{r}).$
Proposition 4.2 (Vector-partition expansion of the $B$-term).
Let $k=N(d)+a$ with $1\leq a\leq p-1$. Then
(4.1) $B_{k}[x^{k}]\equiv a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}+\sum_{\begin{subarray}{c}\Pi\vdash d\\ \Pi\neq(d)\end{subarray}}(-1)^{R(\Pi)-1}(R(\Pi)-1)!\frac{d!}{\prod_{\beta\neq 0}r_{\beta}!(\beta!)^{r_{\beta}}}\prod_{\beta\neq 0}\left(\frac{H_{\beta}(y)}{F_{a}(y)}\right)^{r_{\beta}}\right)\pmod{p}$.
Here $(d)$ denotes the one-block partition of $d$.
Proof.
Let $a^{e}=\prod_{n\geq 0}a_{n}^{e_{n}}$ be a monomial occurring in $B_{k}[x^{k}]$. Then
$\sum_{n\geq 0}e_{n}=p^{2},\qquad\sum_{n\geq 0}ne_{n}=k.$
By (2.5),
$\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=c(e)+d(e)-2,$
where $c(e),d(e)\in\mathbf{Z}_{\geq 0}$ measure the carry defects in the additions
$\sum_{n}e_{n}n=k,\qquad\sum_{n}e_{n}=p^{2}.$
Hence a monomial can survive modulo $p$ only if
$c(e)+d(e)=2.$
We first show that, because $k\equiv a\not\equiv 0\pmod{p}$, one must in fact have
$d(e)=2,\qquad c(e)=0.$
Indeed, if $d(e)=0$, then $\sum_{n}S_{p}(e_{n})=1$. Writing $e_{n}=pu_{n}+r_{n}$ with $0\leq r_{n}\leq p-1$, we have
$\sum_{n}r_{n}\equiv\sum_{n}e_{n}=p^{2}\equiv 0\pmod{p}$.
If some $r_{n}$ were nonzero, then $\sum_{n}r_{n}\geq p$, contradicting $\sum_{n}S_{p}(e_{n})=1$. Thus every $e_{n}$ is divisible by $p$. Dividing by $p$ and repeating the same argument shows that every $e_{n}$ is divisible by $p^{2}$, so exactly one $e_{t}$ equals $p^{2}$. Then $k=p^{2}t$, contradicting $p\nmid k$.
If $d(e)=1$, then $\sum_{n}S_{p}(e_{n})=p$. Again write $e_{n}=pu_{n}+r_{n}$ with $0\leq r_{n}\leq p-1$. Since $\sum_{n}e_{n}=p^{2}$, the sum $\sum_{n}r_{n}$ is a multiple of $p$. If some $r_{n}$ were nonzero, then $\sum_{n}r_{n}\geq p$; because
$\sum_{n}S_{p}(e_{n})=\sum_{n}r_{n}+\sum_{n}S_{p}(u_{n})=p,$
this would force $\sum_{n}r_{n}=p$ and $\sum_{n}S_{p}(u_{n})=0$, hence every $u_{n}=0$ and $\sum_{n}e_{n}=p$, impossible. Therefore all $r_{n}$ vanish, so every $e_{n}$ is divisible by $p$. But then $k=\sum_{n}e_{n}n$ is divisible by $p$, again a contradiction. This proves that $d(e)=2$, hence $c(e)=0$.
Because $c(e)=0$, the addition $\sum_{n}e_{n}n=k$ is carry-free in base $p$. Write each positive index uniquely as
$n=N(\beta)+j,\qquad 0\leq j\leq a,$
where $\beta=(\beta_{1},\dots,\beta_{r})$ records the digits in positions $p,p^{2},\dots,p^{r}$. For $\beta\neq 0$ set
$m_{\beta,j}:=e_{N(\beta)+j},\qquad r_{\beta}:=\sum_{j=0}^{a}m_{\beta,j},$
and for the zero vector set
$n_{j}:=e_{j}\qquad(1\leq j\leq a).$
The carry-free condition is equivalent to
$\sum_{\beta\neq 0}r_{\beta}\beta=d,\qquad\sum_{\beta\neq 0}\sum_{j=0}^{a}jm_{\beta,j}+\sum_{j=1}^{a}jn_{j}=a.$
Thus the nonzero vectors occurring form a vector partition
$\Pi=\prod_{\beta\neq 0}\beta^{r_{\beta}}\qquad\text{of }d.$
The only excluded index is $k=N(d)+a$, so when $\Pi=(d)$ the choice $m_{d,a}=1$ and all other positive multiplicities zero must be removed.
Fix such a partition $\Pi$, and set
$R:=R(\Pi)=\sum_{\beta\neq 0}r_{\beta},\qquad t:=\sum_{j=1}^{a}n_{j},\qquad M:=R+t.$
Then $e_{0}=p^{2}-M$, and the coefficient of the corresponding monomial is
$B_{k}[x^{k}][a^{e}]$ $=k!\,\frac{(p^{2}-1)!}{(p^{2}-M)!\,\prod_{\beta,j}m_{\beta,j}!\,\prod_{j=1}^{a}n_{j}!}\prod_{\beta,j}\frac{1}{(N(\beta)+j)!^{m_{\beta,j}}}\prod_{j=1}^{a}\frac{1}{j!^{n_{j}}}.$
Modulo $p$,
$\frac{(p^{2}-1)!}{(p^{2}-M)!}=\prod_{u=1}^{M-1}(p^{2}-u)\equiv(-1)^{M-1}(M-1)!.$
Also, because the addition of the positive indices is carry-free, Lucas’ theorem in multinomial form gives
$\frac{k!}{\prod_{\beta,j}(N(\beta)+j)!^{m_{\beta,j}}\prod_{j=1}^{a}j!^{n_{j}}}\equiv\frac{a!\,d!}{\prod_{\beta,j}j!^{m_{\beta,j}}(\beta!)^{m_{\beta,j}}\prod_{j=1}^{a}j!^{n_{j}}}\pmod{p}$.
Hence
$B_{k}[x^{k}][a^{e}]$ $\equiv(-1)^{M-1}(M-1)!\,\frac{a!\,d!}{\prod_{\beta,j}m_{\beta,j}!\,j!^{m_{\beta,j}}(\beta!)^{m_{\beta,j}}\,\prod_{j=1}^{a}n_{j}!\,j!^{n_{j}}}\prod_{\beta,j}a_{N(\beta)+j}^{m_{\beta,j}}\prod_{j=1}^{a}a_{j}^{n_{j}}.$
Now sum first over the zero-vector blocks. Since
$F_{a}(y)-1=\sum_{j=1}^{a}\frac{a_{j}}{j!}y^{j},$
we have the negative-binomial identity
$\frac{1}{F_{a}(y)^{R}}=\sum_{t\geq 0}(-1)^{t}\binom{R+t-1}{t}\bigl(F_{a}(y)-1\bigr)^{t}.$
After multiplying by (R1)!(R-1)!, this exactly reproduces the factor
$(-1)^{t}(R+t-1)!$
coming from the zero-vector blocks. Next, for each fixed nonzero $\beta$,
$\frac{1}{r_{\beta}!}\bigl(H_{\beta}(y)\bigr)^{r_{\beta}}=\sum_{\sum_{j}m_{\beta,j}=r_{\beta}}\prod_{j=0}^{a}\frac{1}{m_{\beta,j}!}\left(\frac{a_{N(\beta)+j}}{j!}y^{j}\right)^{m_{\beta,j}}.$
Collecting all $\beta$ therefore yields the contribution of all monomials with fixed vector partition $\Pi$:
$(-1)^{R-1}(R-1)!\,\frac{d!}{\prod_{\beta\neq 0}r_{\beta}!(\beta!)^{r_{\beta}}}\prod_{\beta\neq 0}\left(\frac{H_{\beta}(y)}{F_{a}(y)}\right)^{r_{\beta}}.$
Finally, when $\Pi=(d)$ one must delete the forbidden monomial corresponding to the excluded index $k=N(d)+a$; this is exactly accomplished by replacing $H_{d}(y)$ with the truncated polynomial
$\widetilde{H}_{d}(y)=\sum_{j=0}^{a-1}\frac{a_{N(d)+j}}{j!}y^{j}.$
Extracting the coefficient of $y^{a}$ now gives (4.1). ∎
5. Low coefficients and a multivariate cumulant collapse for $a\neq 0$
The first block of coefficients admits a particularly simple description modulo $p$.
Proposition 5.1 (The first block modulo $p$).
For $0\leq n\leq p-1$ one has
$a_{n}\equiv(-1)^{n}(n+1)^{n-1}\pmod{p}$,
with the convention that the case $n=0$ gives $a_{0}=1$.
Proof.
Let
$g(x):=\sum_{n\geq 0}\frac{a_{n}}{n!}x^{n},$
so that $f_{\phi}(x)=xg(x)$ and $g(0)=1$. The Böttcher equation becomes
$g(x^{q})=g(x)^{q}\bigl(1+p^{2}xg(x)\bigr),\qquad q=p^{2}.$
Taking formal logarithms gives
(5.1) $\log g(x^{q})=q\log g(x)+\log\bigl(1+p^{2}xg(x)\bigr).$
Fix $1\leq n\leq p-1$. Since $q=p^{2}>n$, the coefficient of $x^{n}$ on the left-hand side of (5.1) is 0. Expanding the second logarithm, we obtain
$\log\bigl(1+p^{2}xg(x)\bigr)=p^{2}xg(x)+\sum_{m\geq 2}(-1)^{m-1}\frac{p^{2m}}{m}x^{m}g(x)^{m}.$
Hence the $x^{n}$-coefficient of the second term is congruent modulo $p^{3}$ to the $x^{n}$-coefficient of $p^{2}xg(x)$, because every summand with $m\geq 2$ carries the factor $p^{2m}$ with $2m\geq 4$. Dividing the coefficient relation by $p^{2}$ therefore yields
$x^{n}]\log g(x)\equiv-[x^{n-1}]g(x)\pmod{p}$.
Since $g(x)-F_{p-1}(x)=O(x^{p})$, we conclude that
(5.2) $\log F_{p-1}(x)\equiv-xF_{p-1}(x)\pmod{(p,x^{p})}$.
Set
$U(x):=xF_{p-1}(x)\in x+x^{2}\mathbf{F}_{p}[x]/(x^{p+1}).$
Exponentiating (5.2) in $\mathbf{F}_{p}[x]/(x^{p})$ gives
$F_{p-1}(x)\equiv e^{-xF_{p-1}(x)}\pmod{(p,x^{p})}$,
and therefore
$U(x)=xF_{p-1}(x)\equiv xe^{-U(x)}\pmod{(p,x^{p+1})}$.
Equivalently,
$U(x)e^{U(x)}\equiv x\pmod{(p,x^{p+1})}$.
By Lagrange inversion, the unique formal solution of $U=xe^{-U}$ satisfies
$x^{m}]U(x)=\frac{1}{m}[u^{m-1}]e^{-mu}=\frac{(-m)^{m-1}}{m!}\qquad(m\geq 1).$
Thus, modulo $(p,x^{p+1})$,
$U(x)\equiv\sum_{m=1}^{p}\frac{(-m)^{m-1}}{m!}x^{m}.$
Dividing by $x$ gives
$F_{p-1}(x)\equiv 1+\sum_{n=1}^{p-1}\frac{(-(n+1))^{n}}{(n+1)!}x^{n}\equiv 1+\sum_{n=1}^{p-1}(-1)^{n}(n+1)^{n-1}\frac{x^{n}}{n!}\pmod{(p,x^{p})}$.
Comparing coefficients proves the proposition. ∎
Corollary 5.2.
For every $1\leq a\leq p-1$ one has
$\log F_{a}(y)\equiv-yF_{a}(y)\pmod{(p,y^{a+1})}$,
and hence
$a![y^{a}]\log F_{a}(y)\equiv-a\,a_{a-1}\pmod{p}$.
Proof.
Since $F_{a}(y)-F_{p-1}(y)=O(y^{a+1})$, Proposition 5.1 and (5.2) give
$\log F_{a}(y)\equiv\log F_{p-1}(y)\equiv-yF_{p-1}(y)\equiv-yF_{a}(y)\pmod{(p,y^{a+1})}$.
Taking the coefficient of $y^{a}$ yields
$y^{a}]\log F_{a}(y)\equiv-[y^{a-1}]F_{a}(y)=-\frac{a_{a-1}}{(a-1)!}\pmod{p}$,
and multiplying by $a!$ gives the stated formula. ∎
The expansion (4.1) becomes tractable because, after the inductive substitutions, every block polynomial depends only on the total weight of the corresponding digit vector.
Lemma 5.3 (Induction-to-block transfer).
Assume that (1.1) is already known for every index of the form
$N(\beta)+j\qquad(0\leq j\leq a,\ \beta\neq 0,\ |\beta|<s).$
Then for every nonzero digit vector $\beta$ with $|\beta|<s$ one has
$H_{\beta}(y)\equiv(-D)^{|\beta|}F_{a}(y)\pmod{p},\qquad D:=1+y\frac{d}{dy}.$
Consequently, for every digit vector $d$ of weight $s$,
$\widetilde{H}_{d}(y)\equiv(-D)^{s}F_{a}(y)+(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}\pmod{p}$.
Proof.
Fix a nonzero digit vector $\beta$ with $|\beta|<s$. By the induction hypothesis from (1.1), for every $0\leq j\leq a$ we have
$a_{N(\beta)+j}\equiv(-1)^{|\beta|}(j+1)^{|\beta|}a_{j}\pmod{p}$.
Summing over $j$ gives
$H_{\beta}(y)=\sum_{j=0}^{a}\frac{a_{N(\beta)+j}}{j!}y^{j}\equiv\sum_{j=0}^{a}(-1)^{|\beta|}(j+1)^{|\beta|}\frac{a_{j}}{j!}y^{j}=(-D)^{|\beta|}F_{a}(y)\pmod{p}$,
because
$D^{m}F_{a}(y)=\sum_{j=0}^{a}(j+1)^{m}\frac{a_{j}}{j!}y^{j}\qquad(m\geq 0).$
This proves the first assertion.
Now let $d$ be a digit vector of weight $s$. For $0\leq j\leq a-1$, the same induction hypothesis gives
$a_{N(d)+j}\equiv(-1)^{s}(j+1)^{s}a_{j}\pmod{p}$.
Therefore
$\widetilde{H}_{d}(y)\equiv\sum_{j=0}^{a-1}(-1)^{s}(j+1)^{s}\frac{a_{j}}{j!}y^{j}\pmod{p}$.
Comparing with
$D^{s}F_{a}(y)=\sum_{j=0}^{a}(j+1)^{s}\frac{a_{j}}{j!}y^{j},$
we see that the only missing term is the top coefficient $j=a$, namely
$(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}.$
Hence
$\widetilde{H}_{d}(y)\equiv(-D)^{s}F_{a}(y)+(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}\pmod{p}$.
This proves the lemma. ∎
For nonzero digit vectors define
$U_{\beta}(y):=(-1)^{|\beta|}\frac{D^{|\beta|}F_{a}(y)}{F_{a}(y)}.$
Introduce formal variables $t=(t_{1},\dots,t_{r})$ and set
$t^{\beta}:=\prod_{i=1}^{r}t_{i}^{\beta_{i}}\qquad\text{for }\beta=(\beta_{1},\dots,\beta_{r}).$
The associated multivariate moment series is
$M(t):=1+\sum_{\beta\neq 0}U_{\beta}(y)\frac{t^{\beta}}{\beta!}.$
Define the multivariate cumulants $K_{d}(U)$ by
(5.3) $\log M(t)=\sum_{d\neq 0}K_{d}(U)\frac{t^{d}}{d!}.$
Proposition 5.4 (Collapse to one variable).
Let $d=(d_{1},\dots,d_{r})$ be a nonzero digit vector and set $s:=|d|$. Then
$a![y^{a}]K_{d}(U)=(-1)^{s}a![y^{a}]N^{s}\log F_{a}(y),\qquad N:=y\frac{d}{dy}.$
Consequently,
(5.4) $a![y^{a}]K_{d}(U)\equiv(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
Proof.
Because $U_{\beta}(y)$ depends only on $|\beta|$, the multivariate series depends only on
$T:=t_{1}+\cdots+t_{r}.$
Indeed,
$\sum_{|\beta|=m}\frac{t^{\beta}}{\beta!}=\frac{T^{m}}{m!},$
so
$M(t)$ $=1+\sum_{m\geq 1}(-1)^{m}\frac{D^{m}F_{a}(y)}{F_{a}(y)}\frac{T^{m}}{m!}$
$=\frac{\exp(-TD)F_{a}(y)}{F_{a}(y)}$
$=\frac{e^{-T}F_{a}(ye^{-T})}{F_{a}(y)}.$
Here we used $D=1+N$ and the identity $e^{-TN}G(y)=G(ye^{-T})$ for every polynomial $G$. Taking logarithms gives
$\log M(t)=-T+\log F_{a}(ye^{-T})-\log F_{a}(y).$
Expanding the middle term as a Taylor series in the operator $N$ yields
$\log F_{a}(ye^{-T})=\sum_{m\geq 0}(-1)^{m}N^{m}\log F_{a}(y)\frac{T^{m}}{m!}.$
Since the coefficient of $t^{d}/d!$ in a series depending only on $T$ is the coefficient of $T^{|d|}/|d|!$, we obtain
$a![y^{a}]K_{d}(U)=(-1)^{s}a![y^{a}]N^{s}\log F_{a}(y).$
The linear term T-T contributes only when $s=1$, and even then its $y^{a}$-coefficient is 0 because $a\geq 1$.
Finally,
$a![y^{a}]N^{s}\log F_{a}(y)=a^{s}\,a![y^{a}]\log F_{a}(y),$
so Corollary 5.2 gives
$a![y^{a}]K_{d}(U)=(-1)^{s}a^{s}\,a![y^{a}]\log F_{a}(y)\equiv(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
This proves the proposition. ∎
6. Proof of (1.1)
We now prove (1.1). The case $a=0$ is already known from Theorem 2.3, so it remains to treat the residue classes $1\leq a\leq p-1$.
Proof of (1.1).
We argue by outer induction on $a\in\{0,1,\dots,p-1\}$. The case $a=0$ is Theorem 2.3. Fix $1\leq a\leq p-1$ and assume that the theorem is already known for every smaller residue class $0,1,\dots,a-1$.
For this fixed $a$, we argue by inner induction on
$s=d_{1}+\cdots+d_{r}.$
If $s=0$, then $k=a$ and the desired congruence is tautological.
Now let $s\geq 1$, let $d=(d_{1},\dots,d_{r})$ be a digit vector of weight $s$, and write
$k=N(d)+a.$
By Proposition 2.2 and the already established formula in residue class a1a-1 (for $a=1$ this is Theorem 2.3), we have
(6.1) $C_{k}[x^{k}]\equiv a\,a_{k-1}\equiv(-1)^{s}a^{s+1}a_{a-1}\pmod{p}$.
By Proposition 2.1,
(6.2) $A_{k}[x^{k}]\equiv 0\pmod{p}$.
For the $B$-term, set
$V_{\beta}(y):=\frac{H_{\beta}(y)}{F_{a}(y)}\qquad(\beta\neq 0).$
Let $K_{d}(V)$ be defined by the same cumulant formula as in (5.3), with $U_{\beta}$ replaced by $V_{\beta}$. Since the one-block partition contributes exactly $V_{d}(y)=H_{d}(y)/F_{a}(y)$, Proposition 4.2 may be rewritten as
(6.3) $B_{k}[x^{k}]\equiv a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}+K_{d}(V)-V_{d}(y)\right)\pmod{p}$.
Now every block appearing in a proper partition of $d$ has weight strictly smaller than $s$. Therefore Lemma 5.3, applied with the inner induction hypothesis from (1.1), gives
$V_{\beta}(y)=\frac{H_{\beta}(y)}{F_{a}(y)}\equiv(-1)^{|\beta|}\frac{D^{|\beta|}F_{a}(y)}{F_{a}(y)}=U_{\beta}(y)\pmod{p}$
for every such $\beta$. Therefore the proper-partition contribution is unchanged if we replace $V_{\beta}$ by $U_{\beta}$:
(6.4) $a![y^{a}]\bigl(K_{d}(V)-V_{d}(y)\bigr)\equiv a![y^{a}]\bigl(K_{d}(U)-U_{d}(y)\bigr)\pmod{p}$.
On the other hand, Lemma 5.3 gives
$\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}-U_{d}(y)\equiv(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!F_{a}(y)}\pmod{p}$.
Since $F_{a}(y)=1+O(y)$, we get
(6.5) $a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}-U_{d}(y)\right)=(-1)^{s-1}(a+1)^{s}a_{a}.$
Combining (6.3), (6.4), (6.5), and Proposition 5.4, we obtain
(6.6) $B_{k}[x^{k}]\equiv(-1)^{s-1}(a+1)^{s}a_{a}+(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
Finally, substitute (6.2), (6.6), and (6.1) into the recursion (2.1):
$a_{k}$ $\equiv-\left((-1)^{s-1}(a+1)^{s}a_{a}+(-1)^{s+1}a^{s+1}a_{a-1}\right)-(-1)^{s}a^{s+1}a_{a-1}\pmod{p}$
$=(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
This completes both inductions and proves the theorem. ∎
7. Proof of (1.2) and Theorem 1.2
We first promote Proposition 5.1 to the fully explicit form of Theorem $1.$
Proof of (1.2).
Let $k\geq 1$ and write
$k=a+d_{1}p+d_{2}p^{2}+\cdots+d_{r}p^{r},\qquad 0\leq a,d_{i}\leq p-1,$
with $s=S_{p}(k)-a$.
If $a=0$, then part (i) gives
$a_{k}\equiv(-1)^{s}\pmod{p}$.
Since $S_{p}(k)=s$ and $a+1=1$, this is exactly the desired formula.
Assume now that $1\leq a\leq p-1$. By part (i),
$a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
Applying Proposition 5.1 with $n=a$ gives
$a_{a}\equiv(-1)^{a}(a+1)^{a-1}\pmod{p}$.
Substituting this into the previous congruence yields
$a_{k}\equiv(-1)^{s+a}(a+1)^{s+a-1}\pmod{p}$.
Since $s+a=S_{p}(k)$, this is exactly the asserted formula. ∎
We now deduce the original residue-class conjecture of Salerno and Silverman.
Proof of Theorem 1.2.
The first congruence
$a_{pm}\equiv(-1)^{m}\pmod{p}$
is exactly Theorem 2.3.
For the second congruence, write
pm1=(m1)p+(p1).pm-1=(m-1)p+(p-1).
If $m=1$, Proposition 5.1 gives $a_{p-1}\equiv 0\pmod{p}$. If $m\geq 2$, then (1.1) yields
$a_{pm-1}\equiv(-1)^{m-1}p^{m-1}a_{p-1}\equiv 0\pmod{p}$.
Thus $a_{pm-1}\equiv 0\pmod{p}$ for every $m\geq 1$.
For the third congruence, write
pm2=(m1)p+(p2).pm-2=(m-1)p+(p-2).
Then (1.1) gives
$a_{pm-2}\equiv(-1)^{m-1}(p-1)^{m-1}a_{p-2}\equiv a_{p-2}\pmod{p}$.
By Proposition 5.1,
$a_{p-2}\equiv(-1)^{p-2}(p-1)^{p-3}\equiv-1\pmod{p}$.
Therefore
$a_{pm-2}\equiv-1\pmod{p}$
for every $m\geq 1$. ∎