### 06 Apr 2026 v2

# Digit-sum Formulas and Divisible Classes for Böttcher Coordinates of $x^{p^{2}}+p^{r+2}x^{p^{2}+1}$

Abstract.
Let $p$ be an odd prime and, for $r\geq 0$, let
$\phi_{r}(x)=x^{p^{2}}+p^{r+2}x^{p^{2}+1}.$
For $r=0$ we prove a complete digit-sum congruence formula modulo $p$ for the coefficients of the Böttcher coordinate of $\phi_{0}$. As consequences we prove Conjecture 25 of Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$. For $r\geq 1$ we prove on the divisible non-pure class a leading term theorem together with a global digit-weight lower bound. These results yield an unconditional proof of Conjecture 27(a) of Salerno and Silverman.
Key words and phrases:
Böttcher coordinate, superattracting germ, non-archimedean dynamics
2020 Mathematics Subject Classification:
37P35, 37F50, 11S82
1. Introduction
The local study of superattracting germs goes back to Böttcher’s theorem, which provides a canonical local coordinate in which a germ becomes a pure monomial. In complex dynamics this coordinate plays a central role both near a superattracting fixed point and, for polynomials, at infinity, where it encodes the escape-rate function and the external combinatorics of Julia sets and the Mandelbrot set; see, for example, [1, 2, 3, 6, 7]. The same normal form also appears in higher-dimensional holomorphic dynamics [5].
In the non-archimedean setting, Böttcher coordinates have an additional arithmetic aspect. Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs and formulated several precise conjectures in the wild case [8]. More recently, Fu and Nie proved the expected radius of convergence in a general wild superattracting setting [4]; in particular, their theorem settles Conjecture 27(b) for the family considered here. The family relevant for the present paper is
$\phi_{r}(x)=x^{p^{2}}+p^{r+2}x^{p^{2}+1}\qquad(r\geq 0),$
where $p\geq 3$ is an odd prime. If
$f_{r}(x)=x\sum_{k\geq 0}\frac{a_{k}(r)}{k!}x^{k}$
denotes the Böttcher coordinate characterized by
$\phi_{r}\bigl(f_{r}(x)\bigr)=f_{r}(x^{p^{2}}),$
then Conjectures 25 and 27 of [8] predict striking congruence and valuation patterns for the coefficients $a_{k}(r)$. Since Conjecture 27(b) is already supplied by Fu and Nie, the higher-fiber part of the present paper is concerned only with Conjecture 27(a).
The special fiber $r=0$ already exhibits a rich mod-$p$ structure. In that case one expects a digit-sum formula for every coefficient, not only on the divisible class $p\mid k$ but on every residue class modulo $p$. Our first result establishes the full congruence law.
Theorem 1.1 (Digit-sum formula in the special fiber).
Let $k\geq 1$ and write
$k=a+d_{1}p+d_{2}p^{2}+\cdots+d_{r}p^{r},\qquad 0\leq a,d_{i}\leq p-1,$
with
$s:=d_{1}+\cdots+d_{r}.$
Then
(1.1) $a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
In fact, one has the stronger closed formula
(1.2) $a_{k}\equiv(-1)^{a+s}(a+1)^{a+s-1}\pmod{p}$.
As an immediate consequence we recover the residue-class congruences conjectured in [8, Conjecture 25].
Theorem 1.2 (Salerno–Silverman’s Conjecture 25 for $r=0$).
For every integer $m\geq 1$ one has
$a_{pm}\equiv(-1)^{m},\qquad a_{pm-1}\equiv 0,\qquad a_{pm-2}\equiv-1\pmod{p}$.
For the higher fibers $r\geq 1$, the relevant structure is valuation-theoretic. If
$v_{i}(r):=\operatorname{ord}_{p}\bigl(a_{p^{i}}(r)\bigr)\qquad(i\geq 0),$
and if
$m=m_{0}+m_{1}p+\cdots+m_{s}p^{s},\qquad 0\leq m_{i}\leq p-1,$
we set
$\Lambda_{m}(r):=\sum_{i=0}^{s}m_{i}v_{i+1}(r),\qquad\epsilon_{m}:=m_{0}+\cdots+m_{s}=S_{p}(m).$
Our third main theorem gives the precise leading term on the divisible non-pure class and, from it, proves the valuation statement Conjecture 27(a) of [8].
Theorem 1.3 (Divisible classes and Conjecture 27(a) for $r\geq 1$).
Fix $r\geq 1$. Then the following hold.
  1. (a)
    If $m\geq 1$ is not a power of $p$, then
    $a_{pm}(r)=(-1)^{\epsilon_{m}}p^{\Lambda_{m}(r)}+O\bigl(p^{\Lambda_{m}(r)+1}\bigr),$
    where
    $v_{0}(r)=r,\qquad v_{1}(r)=pr,$
    and, for every $n\geq 2$,
    $v_{n}(r)=\begin{cases}(p+1)p^{n-2}-2+v_{n-2}(r),&\text{if $n$ is even and $r\geq n-1$,}\\[3.0pt] p\,v_{n-1}(r),&\text{otherwise.}\end{cases}$
    In particular,
    $\operatorname{ord}_{p}\bigl(a_{pm}(r)\bigr)=\Lambda_{m}(r).$
  2. (b)
    If $p\mid k$, then
    $\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\frac{1-p^{-r}}{p-1}k+O(1).$
The special-fiber part of the paper is driven by a complete analysis of the AABBCC recursion modulo $p$, together with a vector-partition description of the surviving $B$-term monomials. The higher-fiber part is organized around a simultaneous induction on three main assertions—a global lower bound, a leading theorem on the divisible non-pure class, and a pure-power leading theorem—together with the resulting subadditivity of the digit weight. Once these ingredients are in place, the pure-power branch pattern and Conjecture 27(a) follow.
We do not treat the case $p=2$ here. The arguments for odd $p$ already exhibit the full mechanism, and keeping $p\geq 3$ avoids a number of low-characteristic notational distractions.
2. The recursion and two structural lemmas
In this part we work in the special fiber $r=0$, so $\phi=\phi_{0}$ and $a_{k}=a_{k}(0)$.
Let $q:=p^{2}$. As in [8, §5], write
$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k},\qquad a_{0}=1,$
so that
$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{q}).$
The coefficient comparison may be organized in the form
(2.1) $a_{k}=A_{k}[x^{k}]-B_{k}[x^{k}]-C_{k}[x^{k}],$
where
$A_{k}:=\frac{k!}{q}\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{q\ell},\qquad B_{k}:=\frac{k!}{q}\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q},\qquad C_{k}:=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q+1}.$
Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. For a nonnegative integer
$N=\sum_{i\geq 0}N_{i}p^{i},\qquad 0\leq N_{i}\leq p-1,$
we write
$S_{p}(N):=\sum_{i\geq 0}N_{i},\qquad T_{p}(N):=\frac{N!}{p^{\operatorname{ord}_{p}(N!)}},$
so $T_{p}(N)$ is the prime-to-$p$ part of $N!$. Throughout we use Legendre’s formula
(2.2) $\operatorname{ord}_{p}(n!)=\frac{n-S_{p}(n)}{p-1}$
and the standard digit-sum subadditivity
(2.3) $S_{p}(u+v)\leq S_{p}(u)+S_{p}(v)$
for nonnegative integers $u,v$, together with its obvious extension to finite sums.
Proposition 2.1 (The $A$-term always vanishes modulo $p$).
For every $k\geq 1$ one has
$A_{k}[x^{k}]\equiv 0\pmod{p}$.
Proof.
If $q\nmid k$, then $A_{k}[x^{k}]=0$ because every monomial occurring in $A_{k}$ has degree divisible by $q=p^{2}$.
Assume now that $k=qn=p^{2}n$. Then only the term $\ell=n$ contributes, so
$A_{k}[x^{k}]=\frac{(qn)!}{q\,n!}\,a_{n}=\frac{(qn)!}{p^{2}n!}\,a_{n}.$
Since multiplying by $p^{2}$ only shifts the base-$p$ digits, we have $S_{p}(qn)=S_{p}(n)$. Hence
$\operatorname{ord}_{p}\!\left(\frac{(qn)!}{p^{2}n!}\right)$ $=\operatorname{ord}_{p}((qn)!)-2-\operatorname{ord}_{p}(n!)$
$=\frac{qn-S_{p}(qn)}{p-1}-2-\frac{n-S_{p}(n)}{p-1}$
$=\frac{(p^{2}-1)n}{p-1}-2$
$=(p+1)n-2.$
Since $n\geq 1$ and $p\geq 3$, this is at least $1$. Therefore $p\mid A_{k}[x^{k}]$, so $A_{k}[x^{k}]\equiv 0\pmod{p}$. ∎
For the $B$-term, let $a^{e}=\prod_{n\geq 0}a_{n}^{e_{n}}$ be a monomial occurring in $B_{k}[x^{k}]$. Then
$\sum_{n\geq 0}e_{n}=q,\qquad\sum_{n\geq 0}ne_{n}=k,$
and the coefficient of $a^{e}$ in $B_{k}[x^{k}]$ is
$B_{k}[x^{k}][a^{e}]=\frac{k!}{q}\binom{q}{e_{0},e_{1},\dots}\prod_{n\geq 0}\frac{1}{(n!)^{e_{n}}}.$
A direct application of Legendre’s formula yields
(2.4) $(p-1)\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=-2p+1-S_{p}(k)+\sum_{n\geq 0}S_{p}(e_{n})+\sum_{n\geq 0}e_{n}S_{p}(n).$
Equivalently, with
$c(e):=\frac{\sum_{n\geq 0}e_{n}S_{p}(n)-S_{p}(k)}{p-1},\qquad d(e):=\frac{\sum_{n\geq 0}S_{p}(e_{n})-1}{p-1},$
one has
(2.5) $\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=c(e)+d(e)-2.$
The two quantities $c(e)$ and $d(e)$ measure the carry defects in the base-$p$ additions $\sum_{n}e_{n}n=k$ and $\sum_{n}e_{n}=p^{2}$, respectively.
Proposition 2.2 (A global $C$-term lemma).
If $k\equiv a\pmod{p}$ with $1\leq a\leq p-1$, then
$C_{k}[x^{k}]\equiv a\,a_{k-1}\pmod{p}$.
If $p\mid k$, then
$C_{k}[x^{k}]\equiv 0\pmod{p}$.
Proof.
Let $a^{e}=\prod_{n=0}^{k-1}a_{n}^{e_{n}}$ be a monomial occurring in $C_{k}[x^{k}]$. Since
$C_{k}=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{p^{2}+1},$
its monomials satisfy
$\sum_{n=0}^{k-1}e_{n}=p^{2}+1,\qquad\sum_{n=0}^{k-1}ne_{n}=k-1.$
The coefficient of $a^{e}$ in $C_{k}[x^{k}]$ is
$C_{k}[x^{k}][a^{e}]=k!\binom{p^{2}+1}{e_{0},e_{1},\dots,e_{k-1}}\prod_{n=0}^{k-1}\frac{1}{(n!)^{e_{n}}},$
and a direct Legendre computation gives
(2.6) $(p-1)\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)=-S_{p}(k)-1+\Sigma(e)+T(e),$
where
$\Sigma(e):=\sum_{n=0}^{k-1}S_{p}(e_{n}),\qquad T(e):=\sum_{n=0}^{k-1}e_{n}S_{p}(n).$
Assume first that $p\mid k$. Write $k=p^{\nu}u$ with $\nu\geq 1$ and $p\nmid u$. Subtracting $1$ from $k$ changes the final $\nu$ zeros in the base-$p$ expansion into p1p-1’s and lowers the next digit by $1$, so
$S_{p}(k-1)=S_{p}(k)-1+\nu(p-1).$
Also,
$\Sigma(e)\geq S_{p}(p^{2}+1)=2,\qquad T(e)\geq S_{p}(k-1).$
Therefore every monomial coefficient in $C_{k}[x^{k}]$ satisfies
$(p-1)\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)$ $\geq-S_{p}(k)-1+2+S_{p}(k-1)$
$=\nu(p-1)\geq p-1.$
Hence $\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)\geq 1$ for every monomial, and so $C_{k}[x^{k}]\equiv 0\pmod{p}$.
Now assume that $k\equiv a\pmod{p}$ with $1\leq a\leq p-1$. Then
$S_{p}(k-1)=S_{p}(k)-1.$
Since again $\Sigma(e)\geq 2$ and $T(e)\geq S_{p}(k-1)$, every monomial that survives modulo $p$ must satisfy
(2.7) $\Sigma(e)=2,\qquad T(e)=S_{p}(k-1).$
The condition $\Sigma(e)=2$ leaves exactly two multiplicity patterns:
$e_{t}=p^{2}+1\qquad\text{or}\qquad e_{r}=p^{2},\ e_{s}=1,$
with all other multiplicities equal to 0.
In the first case,
$k-1=(p^{2}+1)t,\qquad T(e)=(p^{2}+1)S_{p}(t).$
If $t>0$, then
$S_{p}(k-1)=S_{p}((p^{2}+1)t)\leq S_{p}(p^{2}t)+S_{p}(t)=2S_{p}(t)<(p^{2}+1)S_{p}(t)=T(e),$
contradicting (2.7). Hence $t=0$, so $k=1$, and the claimed congruence is immediate.
In the second case,
$k-1=p^{2}r+s,\qquad T(e)=p^{2}S_{p}(r)+S_{p}(s).$
If $r>0$, then
$S_{p}(k-1)=S_{p}(p^{2}r+s)\leq S_{p}(r)+S_{p}(s)<p^{2}S_{p}(r)+S_{p}(s)=T(e),$
again a contradiction. Therefore $r=0$, so s=k1s=k-1 and the unique surviving monomial is
$e_{0}=p^{2},\qquad e_{k-1}=1.$
Its coefficient is
$k!\binom{p^{2}+1}{p^{2},1}\frac{1}{(k-1)!}=k(p^{2}+1)\equiv k\equiv a\pmod{p}$.
Thus
$C_{k}[x^{k}]\equiv a\,a_{k-1}\pmod{p}$,
as claimed. ∎
Theorem 2.3 (The case $a=0$).
For every integer $M\geq 1$ one has
$a_{pM}\equiv(-1)^{M}\pmod{p}$.
The proof of Theorem 2.3 is given in the next section.
3. The Case $a=0$
This section gives a complete proof of Theorem 2.3. The main new input is an explicit coefficient identity for a truncated exponential series.
Lemma 3.1 (A truncated-exponential coefficient identity).
For $k\geq 1$, define
$S_{k-1}(y):=\sum_{\ell=0}^{k-1}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
Then
(3.1) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=[y^{k}]e^{-qy/p}-q\frac{(-1)^{k}}{p^{k}k!}.$
Equivalently,
(3.2) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=\frac{(-q/p)^{k}}{k!}-q\frac{(-1)^{k}}{p^{k}k!}.$
Since $q=p^{2}$, this becomes
(3.3) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=\frac{(-p)^{k}}{k!}-q\frac{(-1)^{k}}{p^{k}k!}.$
Proof.
From the exponential series,
$e^{-y/p}=\sum_{\ell=0}^{\infty}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
Hence
$S_{k-1}(y)=e^{-y/p}-R_{k}(y),\qquad R_{k}(y):=\sum_{\ell=k}^{\infty}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
In particular,
$R_{k}(y)=\frac{(-1)^{k}}{p^{k}k!}y^{k}+O(y^{k+1}),$
so
$S_{k-1}(y)=E(y)+U(y),\qquad E(y):=e^{-y/p},\qquad U(y):=-\frac{(-1)^{k}}{p^{k}k!}y^{k}+O(y^{k+1}).$
Now expand the $q$th power:
$(E+U)^{q}=E^{q}+qE^{q-1}U+\sum_{j=2}^{q}\binom{q}{j}E^{q-j}U^{j}.$
Since $\operatorname{ord}_{y}(U)\geq k$, every term with $j\geq 2$ has $y$-adic order at least $2k>k$, hence contributes nothing to the coefficient of $y^{k}$. Therefore
$y^{k}](E+U)^{q}=[y^{k}]E^{q}+[y^{k}](qE^{q-1}U).$
Because $E(0)=1$, we have $E^{q-1}(y)=1+O(y)$, and hence
$y^{k}](E^{q-1}U)=[y^{k}]U=-\frac{(-1)^{k}}{p^{k}k!}.$
Substituting gives
$y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=[y^{k}]E(y)^{q}-q\frac{(-1)^{k}}{p^{k}k!}.$
Finally, $E(y)^{q}=e^{-qy/p}$, which proves (3.1); equation (3.2) follows immediately from the exponential series. ∎
Corollary 3.2.
For every $m\geq 1$,
$y^{m}]\left(\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}\right)^{q}=\frac{p^{m}}{m!}-\frac{q}{p^{m}m!}.$
Proof.
Apply Lemma 3.1 with $k=m$ and replace $y$ by y-y. Since
$\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}=S_{m-1}(-y),$
we obtain
$y^{m}]\left(\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}\right)^{q}=[y^{m}]e^{qy/p}-q\frac{1}{p^{m}m!}.$
Because $q=p^{2}$, we have $e^{qy/p}=e^{py}$, whose $y^{m}$-coefficient is $p^{m}/m!$. ∎
Proposition 3.3 (Structure of the surviving $B$-term monomials).
Assume that $p\mid k$. Write
$B_{k}[x^{k}]=\sum_{e}\gamma_{B}(e)a^{e},$
where the sum is over all multi-indices $e=(e_{0},\dots,e_{k-1})$ satisfying
$\sigma(e):=\sum_{n=0}^{k-1}e_{n}=q,\qquad\nu(e):=\sum_{n=1}^{k-1}ne_{n}=k,\qquad a^{e}:=\prod_{n=0}^{k-1}a_{n}^{e_{n}}.$
Suppose that a coefficient $\gamma_{B}(e)$ has $p$-adic valuation 0. Then either
$a^{e}=a_{0}^{q-p}a_{1}^{p},$
or else every positive index $n$ with $e_{n}>0$ is divisible by $p$.
Proof.
For such $e$, the coefficient of $a^{e}$ in $B_{k}[x^{k}]$ is
$\gamma_{B}(e):=\frac{k!}{q}\binom{q}{e_{0},\dots,e_{k-1}}\prod_{n=0}^{k-1}\frac{1}{(n!)^{e_{n}}}.$
A straightforward calculation with Legendre’s formula gives
(3.4) $(p-1)\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=-2p+1-S_{p}(k)+\sum_{n=0}^{k-1}\bigl(S_{p}(e_{n})+e_{n}S_{p}(n)\bigr).$
It is convenient to write
$c(e):=\frac{\sum_{n=1}^{k-1}e_{n}S_{p}(n)-S_{p}(k)}{p-1},\qquad d(e):=\frac{\sum_{n=0}^{k-1}S_{p}(e_{n})-1}{p-1}.$
By digit-sum subadditivity, both numerators are nonnegative; since $S_{p}(N)\equiv N\pmod{p-1}$ for every integer $N$, both numerators are divisible by p1p-1, so $c(e),d(e)\in\mathbf{Z}_{\geq 0}$. Using
$\sum_{n=0}^{k-1}e_{n}=q=p^{2},\qquad S_{p}(q)=1,$
we can rewrite (3.4) as
(3.5) $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)+d(e)-2.$
Now assume that $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=0$.
Step 1: if some occurring positive index is not divisible by $p$, then $c(e)\geq 1$. Suppose there exists $n\geq 1$ with $e_{n}>0$ and $p\nmid n$. In the base-$p$ sum
$\sum_{n=1}^{k-1}e_{n}n=k,$
the units digit on the left receives a nonzero contribution from such an $n$, while the right-hand side has units digit 0 because $p\mid k$. Therefore at least one carry must occur from the units place to the $p$-place. Equivalently, the digit-sum defect
$\sum_{n=1}^{k-1}e_{n}S_{p}(n)-S_{p}(k)$
is positive, so $c(e)\geq 1$.
Step 2: $d(e)\neq 0$. Assume for contradiction that $d(e)=0$. Then
$\sum_{n=0}^{k-1}S_{p}(e_{n})=1.$
Since every $e_{n}\geq 0$ and $\sum e_{n}=q=p^{2}$, the only possibility is that one index $m$ satisfies $e_{m}=p^{2}$ and all other $e_{n}$ vanish. In that case,
$k=\sum_{n=1}^{k-1}ne_{n}=mp^{2},$
and therefore
$c(e)=\frac{p^{2}S_{p}(m)-S_{p}(p^{2}m)}{p-1}=\frac{(p^{2}-1)S_{p}(m)}{p-1}=(p+1)S_{p}(m)\geq p+1.$
Hence (3.5) gives
$\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)-2\geq p-1>0,$
contradicting the assumption that the valuation is 0. Thus $d(e)\geq 1$.
Step 3: from $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=0$ we obtain $c(e)=d(e)=1$. By Step 1, whenever some occurring positive index is not divisible by $p$ we have $c(e)\geq 1$, and by Step 2 we always have $d(e)\geq 1$. Since (3.5) says
$0=\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)+d(e)-2,$
we must have
$c(e)=d(e)=1.$
Step 4: $d(e)=1$ forces every $e_{n}$ to be divisible by $p$. The condition $d(e)=1$ means that
$\sum_{n=0}^{k-1}S_{p}(e_{n})-S_{p}\!\left(\sum_{n=0}^{k-1}e_{n}\right)=p-1,$
so the addition of the integers $e_{0},e_{1},\dots,e_{k-1}$ to obtain $q=p^{2}$ produces exactly one carry in base $p$. Now suppose some $e_{n}$ has nonzero units digit. Then in the sum
$\sum_{n=0}^{k-1}e_{n}=p^{2},$
there is a carry from the units place to the $p$-place. But the final result $p^{2}$ has base-$p$ digits $(0,0,1)$, so in order for the $p$-place digit also to end up equal to 0, there must be a second carry from the $p$-place to the $p^{2}$-place. This contradicts the fact that there is only one carry. Hence every $e_{n}$ is divisible by $p$.
So we may write
$e_{n}=pf_{n},\qquad\sum_{n=0}^{k-1}f_{n}=p.$
Then
$k=\sum_{n=1}^{k-1}ne_{n}=p\sum_{n=1}^{k-1}f_{n}n,$
so
$S_{p}(k)=S_{p}\!\left(\sum_{n=1}^{k-1}f_{n}n\right)\leq\sum_{n=1}^{k-1}f_{n}S_{p}(n).$
Therefore
$c(e)=\frac{p\sum_{n=1}^{k-1}f_{n}S_{p}(n)-S_{p}(k)}{p-1}\geq\sum_{n=1}^{k-1}f_{n}S_{p}(n).$
Since $c(e)=1$, it follows that
$\sum_{n=1}^{k-1}f_{n}S_{p}(n)\leq 1.$
The left-hand side is a sum of nonnegative integers, and it is positive because $k>0$; thus exactly one positive index $n$ satisfies $f_{n}=1$, and for that index we have $S_{p}(n)=1$. Therefore $n$ is a power of $p$, and the positive part of the monomial is $a_{n}^{p}$.
If this $n$ is not divisible by $p$, then the only possibility is $n=1$, and since $\sum f_{n}=p$ we must have $f_{0}=p-1$, that is,
$e_{0}=p(p-1)=q-p,\qquad e_{1}=p.$
This gives the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$. In all other cases, every positive index that occurs is divisible by $p$. ∎
Corollary 3.4 (Dependence only on the earlier $a_{pi}$).
If $p\mid k$ and $k>p$, then modulo $p$ the quantity $a_{k}$ depends only on the earlier coefficients $a_{pi}$ with $pi<k$.
Proof.
By Propositions 2.1 and 2.2,
$a_{k}\equiv-B_{k}[x^{k}]\pmod{p}\qquad(p\mid k).$
Now Proposition 3.3 shows that any monomial in $B_{k}[x^{k}]$ whose coefficient can survive modulo $p$ is either the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$ or else involves only indices divisible by $p$. If $k>p$, then the exceptional monomial cannot occur, because its weight is
$0\cdot(q-p)+1\cdot p=p,$
not $k$. Therefore every surviving monomial in $B_{k}[x^{k}]$ involves only $a_{pi}$ with $pi<k$. ∎
Lemma 3.5.
The first nontrivial coefficient of $f_{\phi}$ is
$a_{1}=-1.$
Proof.
Write
$f_{\phi}(x)=x+a_{1}x^{2}+O(x^{3}).$
Then
$\phi(f_{\phi}(x))=f_{\phi}(x)^{q}+p^{2}f_{\phi}(x)^{q+1}.$
Expanding to order $x^{q+1}$ gives
$f_{\phi}(x)^{q}=x^{q}+qa_{1}x^{q+1}+O(x^{q+2}),\qquad p^{2}f_{\phi}(x)^{q+1}=p^{2}x^{q+1}+O(x^{q+2}).$
On the other hand,
$f_{\phi}(x^{q})=x^{q}+O(x^{2q}),$
so comparison of the $x^{q+1}$-coefficients in the Böttcher equation yields
$qa_{1}+p^{2}=0.$
Since $q=p^{2}$, we obtain $a_{1}=-1$. ∎
Lemma 3.6.
One has
$a_{p}\equiv-1\pmod{p}$.
Proof.
By Propositions 2.1 and 2.2,
$a_{p}\equiv-B_{p}[x^{p}]\pmod{p}$.
Since there are no positive indices less than $p$ that are divisible by $p$, Proposition 3.3 shows that the only monomial in $B_{p}[x^{p}]$ whose coefficient can be a $p$-adic unit is the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$. Its coefficient is
$\gamma_{p}=\frac{p!}{q}\binom{q}{q-p,p}=\frac{1}{p^{2}}\prod_{j=0}^{p-1}(p^{2}-j).$
Reducing modulo $p$ gives
$\gamma_{p}\equiv\prod_{j=1}^{p-1}(-j)=(-1)^{p-1}(p-1)!\equiv-1\pmod{p}$
by Wilson’s theorem. Using $a_{0}=1$ and Lemma 3.5, we find
$B_{p}[x^{p}]\equiv\gamma_{p}a_{1}^{p}\equiv(-1)(-1)^{p}\equiv 1\pmod{p}$.
Therefore
$a_{p}\equiv-1\pmod{p}$.
Proof of Theorem 2.3.
We argue by induction on $M$. The case $M=1$ is Lemma 3.6.
Assume now that $M\geq 2$ and that
(3.6) $a_{pi}\equiv(-1)^{i}\pmod{p}\qquad(1\leq i<M).$
Applying Corollary 3.4 with $k=pM$, we obtain
(3.7) $a_{pM}\equiv-B_{pM}[x^{pM}]\pmod{p}$,
and only the indices divisible by $p$ can contribute. Therefore, after setting $y=x^{p}$,
(3.8) $B_{pM}[x^{pM}]\equiv\frac{(pM)!}{q}[y^{M}]\left(\sum_{i=0}^{M-1}\frac{a_{pi}}{(pi)!}y^{i}\right)^{q}\pmod{p}$.
Here $a_{0}=1$ corresponds to $i=0$.
For $i\geq 0$, define
$\nu_{i}:=\frac{(pi)!}{p^{i}i!}.$
Then $\nu_{i}\in\mathbf{Z}$ and
(3.9) $\nu_{i}=\prod_{r=1}^{i}\prod_{t=1}^{p-1}\bigl(p(r-1)+t\bigr)\equiv((p-1)!)^{i}\equiv(-1)^{i}\pmod{p}$
by Wilson’s theorem. Hence, by the induction hypothesis (3.6),
$\omega_{i}:=a_{pi}\nu_{i}^{-1}\in\mathbf{Z}_{(p)}$
satisfies $\omega_{i}\equiv 1\pmod{p}$ for $0\leq i<M$ (with $\omega_{0}=1$). Consequently,
(3.10) $\frac{a_{pi}}{(pi)!}=\frac{\omega_{i}}{p^{i}i!}\qquad(0\leq i<M).$
Write $\omega_{i}=1+pz_{i}$ with $z_{i}\in\mathbf{Z}_{(p)}$, and define
$H_{M}(y):=\sum_{i=0}^{M-1}\frac{y^{i}}{p^{i}i!},\qquad L_{M}(y):=\sum_{i=0}^{M-1}z_{i}\frac{y^{i}}{p^{i}i!}.$
Then by (3.10),
$\sum_{i=0}^{M-1}\frac{a_{pi}}{(pi)!}y^{i}=H_{M}(y)+pL_{M}(y).$
We claim that
(3.11) $\frac{p^{M}M!}{q}[y^{M}]\bigl(H_{M}(y)+pL_{M}(y)\bigr)^{q}\equiv\frac{p^{M}M!}{q}[y^{M}]H_{M}(y)^{q}\pmod{p}$.
Indeed,
$\bigl(H_{M}+pL_{M}\bigr)^{q}-H_{M}^{q}=\sum_{j=1}^{q}\binom{q}{j}p^{j}H_{M}^{q-j}L_{M}^{j}.$
Any term contributing to the coefficient of $y^{M}$ in $H_{M}^{q-j}L_{M}^{j}$ has total denominator $p^{M}\prod i!^{e_{i}}$, so
$y^{M}]H_{M}^{q-j}L_{M}^{j}\in p^{-M}\mathbf{Z}_{(p)}.$
Therefore it is enough to show that
$\operatorname{ord}_{p}\!\left(\frac{p^{j}}{q}\binom{q}{j}\right)\geq 1\qquad(1\leq j\leq q).$
Using
$\binom{q}{j}=\frac{q}{j}\binom{q-1}{j-1}$,
we get
$\frac{p^{j}}{q}\binom{q}{j}=\frac{p^{j}}{j}\binom{q-1}{j-1}$,
so
$\operatorname{ord}_{p}\!\left(\frac{p^{j}}{q}\binom{q}{j}\right)\geq j-\operatorname{ord}_{p}(j)\geq 1.$
This proves (3.11).
Now combine (3.7), (3.8), and (3.11). Since $(pM)!=p^{M}M!\nu_{M}$, we obtain
(3.12) $a_{pM}\equiv-\nu_{M}\frac{p^{M}M!}{q}[y^{M}]H_{M}(y)^{q}\pmod{p}$.
By Corollary 3.2,
$y^{M}]H_{M}(y)^{q}=\frac{p^{M}}{M!}-\frac{q}{p^{M}M!}.$
Substituting into (3.12) gives
$a_{pM}\equiv-\nu_{M}\frac{p^{M}M!}{q}\left(\frac{p^{M}}{M!}-\frac{q}{p^{M}M!}\right)\equiv-\nu_{M}\bigl(p^{2M-2}-1\bigr)\pmod{p}$.
Since $M\geq 2$, we have $p^{2M-2}\equiv 0\pmod{p}$, so
$a_{pM}\equiv\nu_{M}\pmod{p}$.
Finally, (3.9) gives $\nu_{M}\equiv(-1)^{M}\pmod{p}$, and hence
$a_{pM}\equiv(-1)^{M}\pmod{p}$.
This completes the induction. ∎
4. Vector partitions and the $B$-term for $a\neq 0$
The higher base-$p$ digits of the index will be encoded by vectors.
Definition 4.1.
Fix $r\geq 1$. A digit vector is an element
$d=(d_{1},\dots,d_{r})\in\{0,1,\dots,p-1\}^{r}.$
Its weight is
$|d|:=d_{1}+\cdots+d_{r},$
and its numeric value is
$N(d):=\sum_{i=1}^{r}d_{i}p^{i}.$
A vector partition of $d$ is a formal product
$\Pi=\prod_{\beta\neq 0}\beta^{r_{\beta}},$
with $r_{\beta}\in\mathbf{Z}_{\geq 0}$ and
$\sum_{\beta\neq 0}r_{\beta}\beta=d.$
The total number of blocks is
$R(\Pi):=\sum_{\beta\neq 0}r_{\beta}.$
Fix $a\in\{0,1,\dots,p-1\}$. For a digit vector $\beta$ define
$H_{\beta}(y):=\sum_{j=0}^{a}\frac{a_{N(\beta)+j}}{j!}y^{j},\qquad F_{a}(y):=\sum_{j=0}^{a}\frac{a_{j}}{j!}y^{j},\qquad\widetilde{H}_{d}(y):=\sum_{j=0}^{a-1}\frac{a_{N(d)+j}}{j!}y^{j}.$
Write also
$d!:=\prod_{i=1}^{r}d_{i}!,\qquad\beta!:=\prod_{i=1}^{r}\beta_{i}!\qquad\text{for }\beta=(\beta_{1},\dots,\beta_{r}).$
Proposition 4.2 (Vector-partition expansion of the $B$-term).
Let $k=N(d)+a$ with $1\leq a\leq p-1$. Then
(4.1) $B_{k}[x^{k}]\equiv a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}+\sum_{\begin{subarray}{c}\Pi\vdash d\\ \Pi\neq(d)\end{subarray}}(-1)^{R(\Pi)-1}(R(\Pi)-1)!\frac{d!}{\prod_{\beta\neq 0}r_{\beta}!(\beta!)^{r_{\beta}}}\prod_{\beta\neq 0}\left(\frac{H_{\beta}(y)}{F_{a}(y)}\right)^{r_{\beta}}\right)\pmod{p}$.
Here $(d)$ denotes the one-block partition of $d$.
Proof.
Let $a^{e}=\prod_{n\geq 0}a_{n}^{e_{n}}$ be a monomial occurring in $B_{k}[x^{k}]$. Then
$\sum_{n\geq 0}e_{n}=p^{2},\qquad\sum_{n\geq 0}ne_{n}=k.$
By (2.5),
$\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=c(e)+d(e)-2,$
where $c(e),d(e)\in\mathbf{Z}_{\geq 0}$ measure the carry defects in the additions
$\sum_{n}e_{n}n=k,\qquad\sum_{n}e_{n}=p^{2}.$
Hence a monomial can survive modulo $p$ only if
$c(e)+d(e)=2.$
We first show that, because $k\equiv a\not\equiv 0\pmod{p}$, one must in fact have
$d(e)=2,\qquad c(e)=0.$
Indeed, if $d(e)=0$, then $\sum_{n}S_{p}(e_{n})=1$. Writing $e_{n}=pu_{n}+r_{n}$ with $0\leq r_{n}\leq p-1$, we have
$\sum_{n}r_{n}\equiv\sum_{n}e_{n}=p^{2}\equiv 0\pmod{p}$.
If some $r_{n}$ were nonzero, then $\sum_{n}r_{n}\geq p$, contradicting $\sum_{n}S_{p}(e_{n})=1$. Thus every $e_{n}$ is divisible by $p$. Dividing by $p$ and repeating the same argument shows that every $e_{n}$ is divisible by $p^{2}$, so exactly one $e_{t}$ equals $p^{2}$. Then $k=p^{2}t$, contradicting $p\nmid k$.
If $d(e)=1$, then $\sum_{n}S_{p}(e_{n})=p$. Again write $e_{n}=pu_{n}+r_{n}$ with $0\leq r_{n}\leq p-1$. Since $\sum_{n}e_{n}=p^{2}$, the sum $\sum_{n}r_{n}$ is a multiple of $p$. If some $r_{n}$ were nonzero, then $\sum_{n}r_{n}\geq p$; because
$\sum_{n}S_{p}(e_{n})=\sum_{n}r_{n}+\sum_{n}S_{p}(u_{n})=p,$
this would force $\sum_{n}r_{n}=p$ and $\sum_{n}S_{p}(u_{n})=0$, hence every $u_{n}=0$ and $\sum_{n}e_{n}=p$, impossible. Therefore all $r_{n}$ vanish, so every $e_{n}$ is divisible by $p$. But then $k=\sum_{n}e_{n}n$ is divisible by $p$, again a contradiction. This proves that $d(e)=2$, hence $c(e)=0$.
Because $c(e)=0$, the addition $\sum_{n}e_{n}n=k$ is carry-free in base $p$. Write each positive index uniquely as
$n=N(\beta)+j,\qquad 0\leq j\leq a,$
where $\beta=(\beta_{1},\dots,\beta_{r})$ records the digits in positions $p,p^{2},\dots,p^{r}$. For $\beta\neq 0$ set
$m_{\beta,j}:=e_{N(\beta)+j},\qquad r_{\beta}:=\sum_{j=0}^{a}m_{\beta,j},$
and for the zero vector set
$n_{j}:=e_{j}\qquad(1\leq j\leq a).$
The carry-free condition is equivalent to
$\sum_{\beta\neq 0}r_{\beta}\beta=d,\qquad\sum_{\beta\neq 0}\sum_{j=0}^{a}jm_{\beta,j}+\sum_{j=1}^{a}jn_{j}=a.$
Thus the nonzero vectors occurring form a vector partition
$\Pi=\prod_{\beta\neq 0}\beta^{r_{\beta}}\qquad\text{of }d.$
The only excluded index is $k=N(d)+a$, so when $\Pi=(d)$ the choice $m_{d,a}=1$ and all other positive multiplicities zero must be removed.
Fix such a partition $\Pi$, and set
$R:=R(\Pi)=\sum_{\beta\neq 0}r_{\beta},\qquad t:=\sum_{j=1}^{a}n_{j},\qquad M:=R+t.$
Then $e_{0}=p^{2}-M$, and the coefficient of the corresponding monomial is
$B_{k}[x^{k}][a^{e}]$ $=k!\,\frac{(p^{2}-1)!}{(p^{2}-M)!\,\prod_{\beta,j}m_{\beta,j}!\,\prod_{j=1}^{a}n_{j}!}\prod_{\beta,j}\frac{1}{(N(\beta)+j)!^{m_{\beta,j}}}\prod_{j=1}^{a}\frac{1}{j!^{n_{j}}}.$
Modulo $p$,
$\frac{(p^{2}-1)!}{(p^{2}-M)!}=\prod_{u=1}^{M-1}(p^{2}-u)\equiv(-1)^{M-1}(M-1)!.$
Also, because the addition of the positive indices is carry-free, Lucas’ theorem in multinomial form gives
$\frac{k!}{\prod_{\beta,j}(N(\beta)+j)!^{m_{\beta,j}}\prod_{j=1}^{a}j!^{n_{j}}}\equiv\frac{a!\,d!}{\prod_{\beta,j}j!^{m_{\beta,j}}(\beta!)^{m_{\beta,j}}\prod_{j=1}^{a}j!^{n_{j}}}\pmod{p}$.
Hence
$B_{k}[x^{k}][a^{e}]$ $\equiv(-1)^{M-1}(M-1)!\,\frac{a!\,d!}{\prod_{\beta,j}m_{\beta,j}!\,j!^{m_{\beta,j}}(\beta!)^{m_{\beta,j}}\,\prod_{j=1}^{a}n_{j}!\,j!^{n_{j}}}\prod_{\beta,j}a_{N(\beta)+j}^{m_{\beta,j}}\prod_{j=1}^{a}a_{j}^{n_{j}}.$
Now sum first over the zero-vector blocks. Since
$F_{a}(y)-1=\sum_{j=1}^{a}\frac{a_{j}}{j!}y^{j},$
we have the negative-binomial identity
$\frac{1}{F_{a}(y)^{R}}=\sum_{t\geq 0}(-1)^{t}\binom{R+t-1}{t}\bigl(F_{a}(y)-1\bigr)^{t}.$
After multiplying by (R1)!(R-1)!, this exactly reproduces the factor
$(-1)^{t}(R+t-1)!$
coming from the zero-vector blocks. Next, for each fixed nonzero $\beta$,
$\frac{1}{r_{\beta}!}\bigl(H_{\beta}(y)\bigr)^{r_{\beta}}=\sum_{\sum_{j}m_{\beta,j}=r_{\beta}}\prod_{j=0}^{a}\frac{1}{m_{\beta,j}!}\left(\frac{a_{N(\beta)+j}}{j!}y^{j}\right)^{m_{\beta,j}}.$
Collecting all $\beta$ therefore yields the contribution of all monomials with fixed vector partition $\Pi$:
$(-1)^{R-1}(R-1)!\,\frac{d!}{\prod_{\beta\neq 0}r_{\beta}!(\beta!)^{r_{\beta}}}\prod_{\beta\neq 0}\left(\frac{H_{\beta}(y)}{F_{a}(y)}\right)^{r_{\beta}}.$
Finally, when $\Pi=(d)$ one must delete the forbidden monomial corresponding to the excluded index $k=N(d)+a$; this is exactly accomplished by replacing $H_{d}(y)$ with the truncated polynomial
$\widetilde{H}_{d}(y)=\sum_{j=0}^{a-1}\frac{a_{N(d)+j}}{j!}y^{j}.$
Extracting the coefficient of $y^{a}$ now gives (4.1). ∎
5. Low coefficients and a multivariate cumulant collapse for $a\neq 0$
The first block of coefficients admits a particularly simple description modulo $p$.
Proposition 5.1 (The first block modulo $p$).
For $0\leq n\leq p-1$ one has
$a_{n}\equiv(-1)^{n}(n+1)^{n-1}\pmod{p}$,
with the convention that the case $n=0$ gives $a_{0}=1$.
Proof.
Let
$g(x):=\sum_{n\geq 0}\frac{a_{n}}{n!}x^{n},$
so that $f_{\phi}(x)=xg(x)$ and $g(0)=1$. The Böttcher equation becomes
$g(x^{q})=g(x)^{q}\bigl(1+p^{2}xg(x)\bigr),\qquad q=p^{2}.$
Taking formal logarithms gives
(5.1) $\log g(x^{q})=q\log g(x)+\log\bigl(1+p^{2}xg(x)\bigr).$
Fix $1\leq n\leq p-1$. Since $q=p^{2}>n$, the coefficient of $x^{n}$ on the left-hand side of (5.1) is 0. Expanding the second logarithm, we obtain
$\log\bigl(1+p^{2}xg(x)\bigr)=p^{2}xg(x)+\sum_{m\geq 2}(-1)^{m-1}\frac{p^{2m}}{m}x^{m}g(x)^{m}.$
Hence the $x^{n}$-coefficient of the second term is congruent modulo $p^{3}$ to the $x^{n}$-coefficient of $p^{2}xg(x)$, because every summand with $m\geq 2$ carries the factor $p^{2m}$ with $2m\geq 4$. Dividing the coefficient relation by $p^{2}$ therefore yields
$x^{n}]\log g(x)\equiv-[x^{n-1}]g(x)\pmod{p}$.
Since $g(x)-F_{p-1}(x)=O(x^{p})$, we conclude that
(5.2) $\log F_{p-1}(x)\equiv-xF_{p-1}(x)\pmod{(p,x^{p})}$.
Set
$U(x):=xF_{p-1}(x)\in x+x^{2}\mathbf{F}_{p}[x]/(x^{p+1}).$
Exponentiating (5.2) in $\mathbf{F}_{p}[x]/(x^{p})$ gives
$F_{p-1}(x)\equiv e^{-xF_{p-1}(x)}\pmod{(p,x^{p})}$,
and therefore
$U(x)=xF_{p-1}(x)\equiv xe^{-U(x)}\pmod{(p,x^{p+1})}$.
Equivalently,
$U(x)e^{U(x)}\equiv x\pmod{(p,x^{p+1})}$.
By Lagrange inversion (see [9, 10]), the unique formal solution of $U=xe^{-U}$ satisfies
$x^{m}]U(x)=\frac{1}{m}[u^{m-1}]e^{-mu}=\frac{(-m)^{m-1}}{m!}\qquad(m\geq 1).$
Thus, modulo $(p,x^{p+1})$,
$U(x)\equiv\sum_{m=1}^{p}\frac{(-m)^{m-1}}{m!}x^{m}.$
Dividing by $x$ gives
$F_{p-1}(x)\equiv 1+\sum_{n=1}^{p-1}\frac{(-(n+1))^{n}}{(n+1)!}x^{n}\equiv 1+\sum_{n=1}^{p-1}(-1)^{n}(n+1)^{n-1}\frac{x^{n}}{n!}\pmod{(p,x^{p})}$.
Comparing coefficients proves the proposition. ∎
Corollary 5.2.
For every $1\leq a\leq p-1$ one has
$\log F_{a}(y)\equiv-yF_{a}(y)\pmod{(p,y^{a+1})}$,
and hence
$a![y^{a}]\log F_{a}(y)\equiv-a\,a_{a-1}\pmod{p}$.
Proof.
Since $F_{a}(y)-F_{p-1}(y)=O(y^{a+1})$, Proposition 5.1 and (5.2) give
$\log F_{a}(y)\equiv\log F_{p-1}(y)\equiv-yF_{p-1}(y)\equiv-yF_{a}(y)\pmod{(p,y^{a+1})}$.
Taking the coefficient of $y^{a}$ yields
$y^{a}]\log F_{a}(y)\equiv-[y^{a-1}]F_{a}(y)=-\frac{a_{a-1}}{(a-1)!}\pmod{p}$,
and multiplying by $a!$ gives the stated formula. ∎
The expansion (4.1) becomes tractable because, after the inductive substitutions, every block polynomial depends only on the total weight of the corresponding digit vector.
Lemma 5.3 (Induction-to-block transfer).
Assume that (1.1) is already known for every index of the form
$N(\beta)+j\qquad(0\leq j\leq a,\ \beta\neq 0,\ |\beta|<s).$
Then for every nonzero digit vector $\beta$ with $|\beta|<s$ one has
$H_{\beta}(y)\equiv(-D)^{|\beta|}F_{a}(y)\pmod{p},\qquad D:=1+y\frac{d}{dy}.$
Consequently, for every digit vector $d$ of weight $s$,
$\widetilde{H}_{d}(y)\equiv(-D)^{s}F_{a}(y)+(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}\pmod{p}$.
Proof.
Fix a nonzero digit vector $\beta$ with $|\beta|<s$. By the induction hypothesis from (1.1), for every $0\leq j\leq a$ we have
$a_{N(\beta)+j}\equiv(-1)^{|\beta|}(j+1)^{|\beta|}a_{j}\pmod{p}$.
Summing over $j$ gives
$H_{\beta}(y)=\sum_{j=0}^{a}\frac{a_{N(\beta)+j}}{j!}y^{j}\equiv\sum_{j=0}^{a}(-1)^{|\beta|}(j+1)^{|\beta|}\frac{a_{j}}{j!}y^{j}=(-D)^{|\beta|}F_{a}(y)\pmod{p}$,
because
$D^{m}F_{a}(y)=\sum_{j=0}^{a}(j+1)^{m}\frac{a_{j}}{j!}y^{j}\qquad(m\geq 0).$
This proves the first assertion.
Now let $d$ be a digit vector of weight $s$. For $0\leq j\leq a-1$, the same induction hypothesis gives
$a_{N(d)+j}\equiv(-1)^{s}(j+1)^{s}a_{j}\pmod{p}$.
Therefore
$\widetilde{H}_{d}(y)\equiv\sum_{j=0}^{a-1}(-1)^{s}(j+1)^{s}\frac{a_{j}}{j!}y^{j}\pmod{p}$.
Comparing with
$D^{s}F_{a}(y)=\sum_{j=0}^{a}(j+1)^{s}\frac{a_{j}}{j!}y^{j},$
we see that the only missing term is the top coefficient $j=a$, namely
$(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}.$
Hence
$\widetilde{H}_{d}(y)\equiv(-D)^{s}F_{a}(y)+(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}\pmod{p}$.
This proves the lemma. ∎
For nonzero digit vectors define
$U_{\beta}(y):=(-1)^{|\beta|}\frac{D^{|\beta|}F_{a}(y)}{F_{a}(y)}.$
Introduce formal variables $t=(t_{1},\dots,t_{r})$ and set
$t^{\beta}:=\prod_{i=1}^{r}t_{i}^{\beta_{i}}\qquad\text{for }\beta=(\beta_{1},\dots,\beta_{r}).$
The associated multivariate moment series is
$M(t):=1+\sum_{\beta\neq 0}U_{\beta}(y)\frac{t^{\beta}}{\beta!}.$
Define the multivariate cumulants $K_{d}(U)$ by
(5.3) $\log M(t)=\sum_{d\neq 0}K_{d}(U)\frac{t^{d}}{d!}.$
Proposition 5.4 (Collapse to one variable).
Let $d=(d_{1},\dots,d_{r})$ be a nonzero digit vector and set $s:=|d|$. Then
$a![y^{a}]K_{d}(U)=(-1)^{s}a![y^{a}]N^{s}\log F_{a}(y),\qquad N:=y\frac{d}{dy}.$
Consequently,
(5.4) $a![y^{a}]K_{d}(U)\equiv(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
Proof.
Because $U_{\beta}(y)$ depends only on $|\beta|$, the multivariate series depends only on
$T:=t_{1}+\cdots+t_{r}.$
Indeed,
$\sum_{|\beta|=m}\frac{t^{\beta}}{\beta!}=\frac{T^{m}}{m!},$
so
$M(t)$ $=1+\sum_{m\geq 1}(-1)^{m}\frac{D^{m}F_{a}(y)}{F_{a}(y)}\frac{T^{m}}{m!}$
$=\frac{\exp(-TD)F_{a}(y)}{F_{a}(y)}$
$=\frac{e^{-T}F_{a}(ye^{-T})}{F_{a}(y)}.$
Here we used $D=1+N$ and the identity $e^{-TN}G(y)=G(ye^{-T})$ for every polynomial $G$. Taking logarithms gives
$\log M(t)=-T+\log F_{a}(ye^{-T})-\log F_{a}(y).$
Expanding the middle term as a Taylor series in the operator $N$ yields
$\log F_{a}(ye^{-T})=\sum_{m\geq 0}(-1)^{m}N^{m}\log F_{a}(y)\frac{T^{m}}{m!}.$
Since the coefficient of $t^{d}/d!$ in a series depending only on $T$ is the coefficient of $T^{|d|}/|d|!$, we obtain
$a![y^{a}]K_{d}(U)=(-1)^{s}a![y^{a}]N^{s}\log F_{a}(y).$
The linear term T-T contributes only when $s=1$, and even then its $y^{a}$-coefficient is 0 because $a\geq 1$.
Finally,
$a![y^{a}]N^{s}\log F_{a}(y)=a^{s}\,a![y^{a}]\log F_{a}(y),$
so Corollary 5.2 gives
$a![y^{a}]K_{d}(U)=(-1)^{s}a^{s}\,a![y^{a}]\log F_{a}(y)\equiv(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
This proves the proposition. ∎
6. Proof of Theorems 1.1 and 1.2
Proof of (1.1)
We now prove (1.1). The case $a=0$ is already known from Theorem 2.3, so it remains to treat the residue classes $1\leq a\leq p-1$.
Proof of (1.1).
We argue by outer induction on $a\in\{0,1,\dots,p-1\}$. The case $a=0$ is Theorem 2.3. Fix $1\leq a\leq p-1$ and assume that the theorem is already known for every smaller residue class $0,1,\dots,a-1$.
For this fixed $a$, we argue by inner induction on
$s=d_{1}+\cdots+d_{r}.$
If $s=0$, then $k=a$ and the desired congruence is tautological.
Now let $s\geq 1$, let $d=(d_{1},\dots,d_{r})$ be a digit vector of weight $s$, and write
$k=N(d)+a.$
By Proposition 2.2 and the already established formula in residue class a1a-1 (for $a=1$ this is Theorem 2.3), we have
(6.1) $C_{k}[x^{k}]\equiv a\,a_{k-1}\equiv(-1)^{s}a^{s+1}a_{a-1}\pmod{p}$.
By Proposition 2.1,
(6.2) $A_{k}[x^{k}]\equiv 0\pmod{p}$.
For the $B$-term, set
$V_{\beta}(y):=\frac{H_{\beta}(y)}{F_{a}(y)}\qquad(\beta\neq 0).$
Let $K_{d}(V)$ be defined by the same cumulant formula as in (5.3), with $U_{\beta}$ replaced by $V_{\beta}$. Since the one-block partition contributes exactly $V_{d}(y)=H_{d}(y)/F_{a}(y)$, Proposition 4.2 may be rewritten as
(6.3) $B_{k}[x^{k}]\equiv a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}+K_{d}(V)-V_{d}(y)\right)\pmod{p}$.
Now every block appearing in a proper partition of $d$ has weight strictly smaller than $s$. Therefore Lemma 5.3, applied with the inner induction hypothesis from (1.1), gives
$V_{\beta}(y)=\frac{H_{\beta}(y)}{F_{a}(y)}\equiv(-1)^{|\beta|}\frac{D^{|\beta|}F_{a}(y)}{F_{a}(y)}=U_{\beta}(y)\pmod{p}$
for every such $\beta$. Therefore the proper-partition contribution is unchanged if we replace $V_{\beta}$ by $U_{\beta}$:
(6.4) $a![y^{a}]\bigl(K_{d}(V)-V_{d}(y)\bigr)\equiv a![y^{a}]\bigl(K_{d}(U)-U_{d}(y)\bigr)\pmod{p}$.
On the other hand, Lemma 5.3 gives
$\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}-U_{d}(y)\equiv(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!F_{a}(y)}\pmod{p}$.
Since $F_{a}(y)=1+O(y)$, we get
(6.5) $a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}-U_{d}(y)\right)=(-1)^{s-1}(a+1)^{s}a_{a}.$
Combining (6.3), (6.4), (6.5), and Proposition 5.4, we obtain
(6.6) $B_{k}[x^{k}]\equiv(-1)^{s-1}(a+1)^{s}a_{a}+(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
Finally, substitute (6.2), (6.6), and (6.1) into the recursion (2.1):
$a_{k}$ $\equiv-\left((-1)^{s-1}(a+1)^{s}a_{a}+(-1)^{s+1}a^{s+1}a_{a-1}\right)-(-1)^{s}a^{s+1}a_{a-1}\pmod{p}$
$=(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
This completes both inductions and proves the theorem. ∎
Proof of (1.2) and Theorem 1.2
We first combine Proposition 5.1 with (1.1) to obtain the closed formula in Theorem $1.$
Proof of (1.2).
Let $k\geq 1$ and write
$k=a+d_{1}p+d_{2}p^{2}+\cdots+d_{r}p^{r},\qquad 0\leq a,d_{i}\leq p-1,$
with $s=S_{p}(k)-a$.
If $a=0$, then (1.1) gives
$a_{k}\equiv(-1)^{s}\pmod{p}$.
Since $S_{p}(k)=s$ and $a+1=1$, this is exactly the desired formula.
Assume now that $1\leq a\leq p-1$. By (1.1),
$a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
Applying Proposition 5.1 with $n=a$ gives
$a_{a}\equiv(-1)^{a}(a+1)^{a-1}\pmod{p}$.
Substituting this into the previous congruence yields
$a_{k}\equiv(-1)^{s+a}(a+1)^{s+a-1}\pmod{p}$.
Since $s+a=S_{p}(k)$, this is exactly the asserted formula. ∎
We now deduce the original residue-class conjecture of Salerno and Silverman.
Proof of Theorem 1.2.
The first congruence
$a_{pm}\equiv(-1)^{m}\pmod{p}$
is exactly Theorem 2.3.
For the second congruence, write
pm1=(m1)p+(p1).pm-1=(m-1)p+(p-1).
If $m=1$, Proposition 5.1 gives $a_{p-1}\equiv 0\pmod{p}$. If $m\geq 2$, then (1.1) yields
$a_{pm-1}\equiv(-1)^{m-1}p^{m-1}a_{p-1}\equiv 0\pmod{p}$.
Thus $a_{pm-1}\equiv 0\pmod{p}$ for every $m\geq 1$.
For the third congruence, write
pm2=(m1)p+(p2).pm-2=(m-1)p+(p-2).
Then (1.1) gives
$a_{pm-2}\equiv(-1)^{m-1}(p-1)^{m-1}a_{p-2}\equiv a_{p-2}\pmod{p}$.
By Proposition 5.1,
$a_{p-2}\equiv(-1)^{p-2}(p-1)^{p-3}\equiv-1\pmod{p}$.
Therefore
$a_{pm-2}\equiv-1\pmod{p}$
for every $m\geq 1$. ∎
7. The family $\phi_{r}$ for $r\geq 1$: recursion and coefficient valuations
Fix a prime $p$ and set $q:=p^{2}$. Let
$\phi_{r}(x)=x^{q}+p^{r+2}x^{q+1}=x^{q}\bigl(1+qp^{r}x\bigr).$
The Böttcher equation
$\phi_{r}\bigl(f_{r}(x)\bigr)=f_{r}(x^{q})$
gives the usual recursion from [8, §4, §7]. Write
$f_{r}(x)=x\sum_{k\geq 0}\frac{a_{k}(r)}{k!}x^{k},\qquad a_{0}(r)=1,$
and define
(7.1) $A_{k}(r):=\frac{k!}{q}\sum_{\ell=0}^{k-1}\frac{a_{\ell}(r)}{\ell!}x^{q\ell},\qquad B_{k}(r):=\frac{k!}{q}\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}(r)}{\ell!}x^{\ell}\right)^{q},\qquad C_{k}(r):=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}(r)}{\ell!}x^{\ell}\right)^{q+1}.$
Then
(7.2) $a_{k}(r)=A_{k}(r)[x^{k}]-B_{k}(r)[x^{k}]-p^{r}C_{k}(r)[x^{k}].$
This is exactly the same formal decomposition as (2.1); only the outer factor $p^{r}$ is new. Since $r$ is fixed throughout the remainder of the paper, we write $A_{k},B_{k},C_{k}$ in place of $A_{k}(r),B_{k}(r),C_{k}(r)$ whenever no confusion can arise.
Recall the Legendre formula
(7.3) $\operatorname{ord}_{p}(N!)=\frac{N-S_{p}(N)}{p-1}$
for the sum-of-digits function $S_{p}(N)$.
The $A$-term
If $q\nmid k$ then $A_{k}[x^{k}]=0$. If $k=qm$, then
(7.4) $A_{k}[x^{k}]=\frac{(qm)!}{q\,m!}a_{m}.$
Since $S_{p}(qm)=S_{p}(m)$, Legendre’s formula gives
(7.5) $\operatorname{ord}_{p}\!\left(\frac{(qm)!}{q\,m!}\right)=(p+1)m-2.$
The $B$- and $C$-terms
The combinatorial coefficient formulas for $B_{k}[x^{k}]$ and $C_{k}[x^{k}]$ are identical to those in Sections 2 and 3, since the parameter $r$ enters only through the outer factor $p^{r}$ in (7.2). Accordingly, the valuation formulas (3.4) and (3.5) remain valid verbatim for the coefficients of $B_{k}[x^{k}]$. In particular, all $B$-coefficients are $p$-integral, and when $p\mid k$ every unit-coefficient monomial is governed by Proposition 3.3: either it is the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$, or else all of its positive indices are divisible by $p$.
For the $C$-term we use the coefficient formula proved in Proposition 2.2. In the present family this gives the same divisibility statement as before: if $p\mid k$, then every coefficient of $C_{k}[x^{k}]$ is divisible by $p$. We record this as
(7.6) $\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)\geq 1\qquad(p\mid k)$
for every monomial $a^{e}$ occurring in $C_{k}[x^{k}]$.
8. The simultaneous induction for $r\geq 1$
For $n\geq 0$ put
$v_{n}(r):=\operatorname{ord}_{p}\bigl(a_{p^{n}}(r)\bigr).$
When $r$ is fixed we often suppress it from the notation and write simply $v_{n}$. Thus $v_{0}(r)=r$ because comparing the coefficient of $x^{q+1}$ in the Böttcher relation gives
(8.1) $a_{1}(r)=-p^{r}.$
Definition 8.1.
If $k=\sum_{i\geq 0}k_{i}p^{i}$ is the base-$p$ expansion of $k$, define its digit weight by
(8.2) $\Lambda_{r}(k):=\sum_{i\geq 0}k_{i}v_{i}(r)$
and its digit monomial by
(8.3) $\mathfrak{m}_{r}(k):=\prod_{i\geq 0}a_{p^{i}}(r)^{k_{i}}.$
Then
$\operatorname{ord}_{p}\bigl(\mathfrak{m}_{r}(k)\bigr)=\Lambda_{r}(k).$
Proposition 8.2 (Exact two-term decomposition on pure powers).
For every $n\geq 2$ there exist polynomials
$R_{n},S_{n}\in\mathbf{Z}_{p}[a_{1}(r),\dots,a_{p^{n}-1}(r)]$
such that
(8.4) $a_{p^{n}}(r)=\alpha_{n}a_{p^{n-2}}(r)-\gamma_{n}a_{p^{n-1}}(r)^{p}+pR_{n}+p^{r+1}S_{n},$
where
(8.5) $\alpha_{n}:=\frac{(p^{n})!}{q\,(p^{n-2})!},\qquad\operatorname{ord}_{p}(\alpha_{n})=(p+1)p^{n-2}-2,\qquad p^{-\operatorname{ord}_{p}(\alpha_{n})}\alpha_{n}\equiv 1\pmod{p}$,
and
(8.6) $\gamma_{n}\in\mathbf{Z}_{p}^{\times},\qquad\gamma_{n}\equiv-1\pmod{p}$.
Proof.
When $k=p^{n}=qp^{n-2}$, only the term $\ell=p^{n-2}$ contributes to $A_{k}[x^{k}]$, so
$A_{k}[x^{k}]=\frac{(p^{n})!}{q\,(p^{n-2})!}a_{p^{n-2}}=\alpha_{n}a_{p^{n-2}}.$
The valuation statement for $\alpha_{n}$ is exactly (7.5). Its unit congruence follows from Legendre’s formula together with the prime-to-$p$ factorial congruence $T_{p}(p^{m})\equiv(-1)^{m}\pmod{p}$ used in [8, Lemma 22(a)].
On the $B$-side, Theorem 5 of [8] shows that in $B_{p^{n}}[x^{p^{n}}]$ the unique monomial whose coefficient is a $p$-adic unit is the exceptional monomial $a_{0}^{q-p}a_{p^{n-1}}^{p}$. Its coefficient is
$\gamma_{n}=\frac{(p^{n})!}{q}\binom{q}{q-p,p}\frac{1}{(p^{n-1}!)^{p}}=\frac{(p^{n})!}{p\,(p^{n-1}!)^{p}}\binom{p^{2}-1}{p-1}$.
Since
$\operatorname{ord}_{p}\!\left(\frac{(p^{n})!}{(p^{n-1}!)^{p}}\right)=1,$
the first factor is a $p$-adic unit. Writing it in terms of prime-to-$p$ parts gives
$\frac{(p^{n})!}{p\,(p^{n-1}!)^{p}}=\frac{T_{p}(p^{n})}{T_{p}(p^{n-1})^{p}}\equiv\frac{(-1)^{n}}{((-1)^{n-1})^{p}}\equiv-1\pmod{p}$,
because $p$ is odd. Moreover Lucas’s theorem gives
$\binom{p^{2}-1}{p-1}\equiv\binom{p-1}{0}\binom{p-1}{p-1}\equiv 1\pmod{p}$.
Hence $\gamma_{n}\equiv-1\pmod{p}$.
All remaining $B$-monomials have coefficients divisible by $p$, and this is exactly the polynomial $pR_{n}$. Finally, because $p\mid p^{n}$, every coefficient of $C_{p^{n}}[x^{p^{n}}]$ is divisible by $p$ by (7.6). Multiplying by the outer factor $p^{r}$ in (7.2) gives the term $p^{r+1}S_{n}$. ∎
We now organize the argument for $r\geq 1$ as a simultaneous strong induction. We consider the following recursive assertions:
Statement 8.3 (Global digit-weight lower bound).
For every positive integer $k$, one has
$\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)\geq\Lambda_{r}(k).$
Statement 8.4 (Leading term on the divisible non-pure class).
If $p\mid k$ and $k$ is not a power of $p$, then
(8.7) $a_{k}(r)\equiv\mathfrak{m}_{r}(k)\pmod{p^{\Lambda_{r}(k)+1}}$.
Statement 8.5 (Pure-power leading theorem).
For every integer $n\geq 2$, define
(8.8) $A_{n}(r):=(p+1)p^{n-2}-2+v_{n-2}(r),\qquad B_{n}(r):=p\,v_{n-1}(r).$
Then
(8.9) $v_{n}(r)=\min\{A_{n}(r),B_{n}(r)\},$
and, whenever $A_{n}(r)\neq B_{n}(r)$,
(8.10) $a_{p^{n}}(r)=\begin{cases}\alpha_{n}a_{p^{n-2}}(r)+O\bigl(p^{v_{n}(r)+1}\bigr),&\text{if }A_{n}(r)<B_{n}(r),\\[3.0pt] -\gamma_{n}a_{p^{n-1}}(r)^{p}+O\bigl(p^{v_{n}(r)+1}\bigr),&\text{if }B_{n}(r)<A_{n}(r).\end{cases}$
In particular,
$v_{n}(r)\leq p\,v_{n-1}(r)\qquad(n\geq 1).$
Statement 8.6 (Subadditivity of the digit weight).
For all $m,n\geq 0$ one has
$\Lambda_{r}(m+n)\leq\Lambda_{r}(m)+\Lambda_{r}(n).$
More generally, for every finite sum $N=n_{1}+\cdots+n_{t}$,
$\Lambda_{r}(N)\leq\Lambda_{r}(n_{1})+\cdots+\Lambda_{r}(n_{t}).$
To keep the induction parameters explicit, we proceed as follows. For each $n\geq 2$ we first prove Statement 8.5 at the pure-power index $p^{n}$ from Statements 8.3 and 8.4 below $p^{n}$. Once Statement 8.5 is known up to level $n$, the inequalities $v_{j}(r)\leq p\,v_{j-1}(r)$ for $1\leq j\leq n$ imply Statement 8.6 for every integer whose base-$p$ expansion uses only the digits $1,p,\dots,p^{n}$. We then prove Statement 8.4 at each divisible non-pure index $k$, and finally Statement 8.3 at the general index $k$.
Proof of Statement 8.5 at level $n$ and Statement 8.6 up to level $n$.
Fix $n\geq 2$. Assume Statement 8.3 and Statement 8.4 hold for every $k<p^{n}$, and Statement 8.5 holds for every level $2\leq j<n$. By the induction scheme explained above, Statement 8.6 is therefore already available for every integer whose highest base-$p$ digit is at most $p^{n-1}$. In particular,
$v_{j}(r)\leq p\,v_{j-1}(r)\qquad(1\leq j\leq n-1),$
because this is part of Statement 8.5 for $j\geq 2$, while for $j=1$ it is simply $v_{1}(r)=pr=pv_{0}(r)$.
We first control the term $pR_{n}$. The polynomial $R_{n}$ is the sum of the non-exceptional monomials occurring in $B_{p^{n}}[x^{p^{n}}]$. If the coefficient of such a monomial is divisible by $p$, then the outer factor $p$ in $pR_{n}$ already contributes one extra $p$. If the coefficient is a $p$-adic unit, then Proposition 3.3 forces every positive index in that monomial to be divisible by $p$. For each lower divisible non-pure index we may then replace $a_{m}(r)$ by its leading digit monomial using Statement 8.4; on the pure-power line the same replacement is tautological from the definition of $\mathfrak{m}_{r}$. After this substitution, the only carry-free projected monomial of weight $B_{n}(r)=p\,v_{n-1}(r)$ is the exceptional monomial $a_{p^{n-1}}(r)^{p}$, and that monomial has been removed from $R_{n}$. Hence every term in $pR_{n}$ has valuation strictly larger than $B_{n}(r)$, and therefore strictly larger than $\min\{A_{n}(r),B_{n}(r)\}$.
For $p^{r+1}S_{n}$, every coefficient of $C_{p^{n}}[x^{p^{n}}]$ is divisible by $p$ by (7.6). Every monomial in $S_{n}$ is a product of lower coefficients whose total index is $p^{n}-1$. Hence Statement 8.3 for lower indices and Statement 8.6 at lower levels show that every monomial contribution to $p^{r+1}S_{n}$ has valuation at least
$r+1+\Lambda_{r}(p^{n}-1).$
Now $\Lambda_{r}(p^{n}-1)=(p-1)\sum_{i=0}^{n-1}v_{i}(r)$, because the base-$p$ expansion of $p^{n}-1$ has all digits equal to p1p-1 below level $n$. Using $v_{j+1}(r)\leq p\,v_{j}(r)$ for $0\leq j\leq n-2$, we obtain
$\sum_{i=0}^{n-2}(p-1)v_{i}(r)\geq\sum_{i=0}^{n-2}\bigl(v_{i+1}(r)-v_{i}(r)\bigr)=v_{n-1}(r)-v_{0}(r).$
Therefore
$r+1+\Lambda_{r}(p^{n}-1)$ $=1+v_{0}(r)+(p-1)\sum_{i=0}^{n-1}v_{i}(r)$
$\geq 1+v_{0}(r)+v_{n-1}(r)-v_{0}(r)+(p-1)v_{n-1}(r)$
$=1+p\,v_{n-1}(r)=1+B_{n}(r).$
Hence every contribution in $p^{r+1}S_{n}$ has valuation strictly larger than $B_{n}(r)$, and therefore strictly larger than $\min\{A_{n}(r),B_{n}(r)\}$.
Substituting these bounds into Proposition 8.2 gives (8.9) and, whenever $A_{n}(r)\neq B_{n}(r)$, the leading branch formula (8.10). The displayed inequality $v_{n}(r)\leq p\,v_{n-1}(r)$ follows immediately from (8.9).
To deduce Statement 8.6 up to level $n$, let $N=n_{1}+\cdots+n_{t}$ be a finite sum whose highest base-$p$ digit occurs at level at most $n$. Passing from the separate base-$p$ expansions of the $n_{i}$ to the base-$p$ expansion of $N$ amounts to replacing carries of $p$ copies of $p^{j}$ by one copy of $p^{j+1}$. Since $v_{j+1}(r)\leq p\,v_{j}(r)$ for every $j\leq n-1$, each carry can only decrease (or preserve) the total weight. Iterating over all carries yields
$\Lambda_{r}(N)\leq\Lambda_{r}(n_{1})+\cdots+\Lambda_{r}(n_{t}),$
which proves Statement 8.6 for all integers supported in the first $n+1$ digit places. ∎
Proof of Statement 8.4 at the index $k=pm$.
Let $k=pm$ with $m\geq 1$ not a power of $p$. Assume Statement 8.3 and Statement 8.4 hold for every $n^{\prime}<k$, and Statement 8.5 holds for every pure-power level $p^{j}<k$. Then Statement 8.6 is available for all integers $\leq k$.
Claim 8.7 (Shift inequality).
For every integer $m^{\prime}\geq 0$ with $p^{2}m^{\prime}\leq k$ one has
$\Lambda_{r}(p^{2}m^{\prime})\leq(p+1)m^{\prime}-2+\Lambda_{r}(m^{\prime}).$
If $m^{\prime}$ is not a power of $p$, then the inequality is strict.
Proof of the claim.
Write $m^{\prime}=\sum_{i}m^{\prime}_{i}p^{i}$. Since Statement 8.5 has already been proved at every level appearing in the base-$p$ expansion of $m^{\prime}$, we have
$v_{i+2}(r)\leq(p+1)p^{i}-2+v_{i}(r)$
for every $i$ with $m^{\prime}_{i}>0$. Therefore
$\Lambda_{r}(p^{2}m^{\prime})=\sum_{i}m^{\prime}_{i}v_{i+2}(r)\leq\sum_{i}m^{\prime}_{i}\bigl((p+1)p^{i}-2+v_{i}(r)\bigr)=(p+1)m^{\prime}-2\sum_{i}m^{\prime}_{i}+\Lambda_{r}(m^{\prime}).$
Since $\sum_{i}m^{\prime}_{i}\geq 1$, this proves the first assertion. If $m^{\prime}$ is not a power of $p$, then $\sum_{i}m^{\prime}_{i}\geq 2$, so the inequality is strict. ∎
If $q\nmid k$, then $A_{k}[x^{k}]=0$. If $k=qm^{\prime}$, then $m^{\prime}$ is not a power of $p$, so Claim 8.7 is strict. Hence
$\operatorname{ord}_{p}\bigl(A_{k}[x^{k}]\bigr)\geq(p+1)m^{\prime}-2+\Lambda_{r}(m^{\prime})>\Lambda_{r}(k).$
For the $C$-term, every monomial coefficient is divisible by $p$ because $p\mid k$; see (7.6). Therefore
$\operatorname{ord}_{p}\bigl(p^{r}C_{k}[x^{k}]\bigr)\geq 1+r+\Lambda_{r}(k-1)>\Lambda_{r}(k)$
by Statement 8.6. Thus
(8.11) $a_{k}(r)\equiv-B_{k}[x^{k}]\pmod{p^{\Lambda_{r}(k)+1}}$.
If a monomial in $B_{k}[x^{k}]$ contains some positive index not divisible by $p$, then either its coefficient is divisible by $p$ or, by Proposition 3.3, it cannot be a unit-coefficient monomial on the class $p\mid k$. In either case Statement 8.3 shows that its total valuation is $\Lambda_{r}(k)$. Hence modulo $p^{\Lambda_{r}(k)+1}$ one may keep only the divisible truncation and write
(8.12) $B_{k}[x^{k}]\equiv\frac{k!}{q}[y^{m}]\left(\sum_{i=0}^{m-1}\frac{a_{pi}(r)}{(pi)!}y^{i}\right)^{q}\pmod{p^{\Lambda_{r}(k)+1}}$.
For every $i<m$, either $pi$ is a pure power, in which case $a_{pi}(r)=\mathfrak{m}_{r}(pi)$ by definition, or $pi$ is divisible and non-pure, in which case the induction hypothesis gives
$a_{pi}(r)\equiv\mathfrak{m}_{r}(pi)\pmod{p^{\Lambda_{r}(pi)+1}}$.
Because the coefficient in front of each monomial in (8.12) is $p$-integral, every occurrence of one of the error terms raises the total valuation by at least one. Therefore
(8.13) $B_{k}[x^{k}]\equiv\frac{k!}{q}[y^{m}]\left(\sum_{i=0}^{m-1}\frac{\mathfrak{m}_{r}(pi)}{(pi)!}y^{i}\right)^{q}\pmod{p^{\Lambda_{r}(k)+1}}$.
Set
$H_{m}(y):=\sum_{i=0}^{m-1}\frac{\mathfrak{m}_{r}(pi)}{(pi)!}y^{i}.$
We now determine the coefficient of the monomial $\mathfrak{m}_{r}(k)$ in $\frac{k!}{q}[y^{m}]H_{m}(y)^{q}$ modulo $p$. After projecting to $\mathfrak{m}_{r}(k)$, only carry-free tails survive. Write
$m=m_{0}+m_{1}p+\cdots+m_{s}p^{s},\qquad 0\leq m_{\nu}\leq p-1.$
Running the coefficient extraction digit by digit, exactly as in the handwritten digit argument, one reduces the scalar coefficient to successive one-digit problems. At each nonterminal digit level one encounters the full truncated exponential coming from the allowable tails, and Corollary 3.2 shows that its normalized contribution is congruent to $1\pmod{p}$. At the terminal level—the highest nonzero digit of mm—one gets the proper truncation, and the same one-digit calculation as in the base case gives contribution $1\pmod{p}$. Therefore the total coefficient of $\mathfrak{m}_{r}(k)$ in $\frac{k!}{q}[y^{m}]H_{m}(y)^{q}$ is congruent to $1\pmod{p}$.
Equivalently,
$B_{k}[x^{k}]\equiv-\mathfrak{m}_{r}(k)\pmod{p^{\Lambda_{r}(k)+1}}$.
Combining this with (8.11) and (8.13) yields
$a_{k}(r)\equiv\mathfrak{m}_{r}(k)\pmod{p^{\Lambda_{r}(k)+1}}$,
which is Statement 8.4 at the index $k$. ∎
Proof of Statement 8.3 at the index $k$.
Fix $k\geq 1$. Assume Statement 8.3 holds for every $n^{\prime}<k$, Statement 8.4 holds for every divisible non-pure $n^{\prime}<k$, and Statement 8.5 holds for every pure-power level $p^{j}\leq k$. Then Statement 8.6 is available for all integers $\leq k$.
If $k=p^{n}$ is a pure power, then Statement 8.5 gives
$\operatorname{ord}_{p}\bigl(a_{p^{n}}(r)\bigr)=v_{n}(r)=\Lambda_{r}(p^{n}),$
so the lower bound holds with equality. We may therefore assume that $k$ is not a pure power.
We use the recursion (7.2).
The AA-term. If $q\nmid k$, then $A_{k}[x^{k}]=0$. If $k=qm$, then by (7.4), the induction hypothesis, and Claim 8.7,
$\operatorname{ord}_{p}\bigl(A_{k}[x^{k}]\bigr)\geq(p+1)m-2+\Lambda_{r}(m)\geq\Lambda_{r}(qm)=\Lambda_{r}(k).$
The $B$-term. Take a monomial $\gamma_{B}(e)a^{e}$ in $B_{k}[x^{k}]$. By (3.5), $\gamma_{B}(e)\in\mathbf{Z}_{p}$. By the induction hypothesis,
$\operatorname{ord}_{p}(a^{e})\geq\sum_{n=0}^{k-1}e_{n}\Lambda_{r}(n).$
Since $\sum_{n=0}^{k-1}e_{n}n=k$, Statement 8.6 gives
$\sum_{n=0}^{k-1}e_{n}\Lambda_{r}(n)\geq\Lambda_{r}(k).$
Hence every $B$-monomial has valuation at least $\Lambda_{r}(k)$, so
$\operatorname{ord}_{p}\bigl(B_{k}[x^{k}]\bigr)\geq\Lambda_{r}(k).$
The $C$-term. Take a monomial $\gamma_{C}(e)a^{e}$ in $C_{k}[x^{k}]$. By the coefficient formula in Proposition 2.2, the coefficient $\gamma_{C}(e)$ is $p$-integral, and by the induction hypothesis,
$\operatorname{ord}_{p}(a^{e})\geq\sum_{n=0}^{k-1}e_{n}\Lambda_{r}(n)\geq\Lambda_{r}(k-1)$
because $\sum_{n=0}^{k-1}e_{n}n=k-1$. Therefore
$\operatorname{ord}_{p}\bigl(p^{r}\gamma_{C}(e)a^{e}\bigr)\geq r+\Lambda_{r}(k-1)\geq\Lambda_{r}(k),$
where the last inequality is the special case k=(k1)+1k=(k-1)+1 of Statement 8.6. Since every summand in (7.2) has valuation at least $\Lambda_{r}(k)$, so does $a_{k}(r)$. ∎
The pure-power branch pattern and Conjecture 27(a)
We now return to the pure powers and deduce the branch pattern and the asymptotic law predicted in Conjecture 27(a) of [8]. For $n\geq 2$, Statement 8.5 reads
$v_{n}(r)=\min\{A_{n}(r),B_{n}(r)\},\qquad A_{n}(r):=(p+1)p^{n-2}-2+v_{n-2}(r),\qquad B_{n}(r):=p\,v_{n-1}(r).$
Again we suppress the argument $(r)$ whenever no confusion can arise, and we set $s:=\left\lfloor\frac{r+1}{2}\right\rfloor$.
Proposition 8.8 (Pure-power branch pattern).
For every $j\geq 1$ one has:
  1. (1)
    every odd layer $n=2j+1$ is $B$-dominated;
  2. (2)
    an even layer $n=2j$ is $A$-dominated if and only if $r\geq 2j-1$;
  3. (3)
    once $n\geq 2s+1$, all subsequent layers are $B$-dominated.
Proof.
The proof is now unconditional because Statement 8.5 has already been proved.
The initial values are
$v_{0}=r,\qquad v_{1}=pr,\qquad v_{2}=r+p-1,\qquad v_{3}=p(r+p-1).$
Assume that $2j$ is $A$-dominated, so $v_{2j}=A_{2j}$. Then the previous odd layer is $B$-dominated, hence $v_{2j-1}=pv_{2j-2}$. Therefore
$A_{2j+1}-B_{2j+1}$ $=\bigl((p+1)p^{2j-1}-2\bigr)+v_{2j-1}-pv_{2j}$
$=\bigl((p+1)p^{2j-1}-2\bigr)+pv_{2j-2}-p\bigl((p+1)p^{2j-2}-2+v_{2j-2}\bigr)$
$=2p-2>0,$
so every odd layer is $B$-dominated.
Now assume that up to level 2j22j-2 the pattern is alternating $B,A,B,A,\dots,B,A$. Then
$v_{2j-2}=r+\sum_{m=1}^{j-1}\bigl((p+1)p^{2m-2}-2\bigr)=r-2(j-1)+\frac{p^{2j-2}-1}{p-1}.$
Hence
$A_{2j}-B_{2j}$ $=\bigl((p+1)p^{2j-2}-2\bigr)+v_{2j-2}-p^{2}v_{2j-2}$
$=-(p-1)\bigl((p+1)(r-2j)+2p+1\bigr).$
Thus $A_{2j}<B_{2j}$ if and only if $r\geq 2j-1$. This proves the second assertion.
Finally, if $n\geq 2s+1$ is odd, then it is $B$-dominated by part (1). If $n\geq 2s+2$ is even, write $n=2j$. Then $j\geq s+1$, so r<2j1r<2j-1 by the definition of $s$, and part (2) shows that $n$ is $B$-dominated. This proves part (3).
Proposition 8.9 (Normalized pure-power units).
For every $n\geq 0$ one has
$p^{-v_{n}(r)}a_{p^{n}}(r)\equiv-1\pmod{p}$.
Proof.
We argue by induction on $n$. The case $n=0$ is (8.1). For $n=1$, one has $A_{p}[x^{p}]=0$, while $p^{r}C_{p}[x^{p}]$ has valuation at least $pr+1$ by (7.6). The unique $B$-monomial of valuation $pr$ is the exceptional monomial $a_{0}^{q-p}a_{1}(r)^{p}$, and its coefficient is the analogous unit $\gamma_{1}$ obtained by the same calculation as in the proof of Proposition 8.2; in particular $\gamma_{1}\equiv-1\pmod{p}$. Hence
$a_{p}(r)\equiv-\gamma_{1}a_{1}(r)^{p}\equiv-p^{pr}\pmod{p^{pr+1}}$,
so the claim holds for $n=1$.
Now let $n\geq 2$ and assume the claim for all smaller indices. By Proposition 8.8, the layer $n$ is either $A$-dominated or $B$-dominated, so Statement 8.5 applies with one of the branches in (8.10). If $A_{n}(r)<B_{n}(r)$, then
$p^{-v_{n}(r)}a_{p^{n}}(r)\equiv p^{-A_{n}(r)}\alpha_{n}a_{p^{n-2}}(r)\equiv-1\pmod{p}$,
because $p^{-\operatorname{ord}_{p}(\alpha_{n})}\alpha_{n}\equiv 1\pmod{p}$ and the induction hypothesis gives the unit of $a_{p^{n-2}}(r)$. If $B_{n}(r)<A_{n}(r)$, then (8.6) and the induction hypothesis yield
$p^{-v_{n}(r)}a_{p^{n}}(r)\equiv p^{-B_{n}(r)}\bigl(-\gamma_{n}a_{p^{n-1}}(r)^{p}\bigr)\equiv(-\gamma_{n})\bigl(p^{-v_{n-1}(r)}a_{p^{n-1}}(r)\bigr)^{p}\equiv 1\cdot(-1)^{p}\equiv-1\pmod{p}$.
Thus the claim holds for every $n\geq 0$. ∎
Corollary 8.10 (Stable pure-power slope).
For every $n\geq 2s$ one has
$v_{n}(r)=\frac{1-p^{-r}}{p-1}p^{n}.$
Proof.
At level $2s$ one computes directly from the closed form in the proof of Proposition 8.8 that
$v_{2s}(r)=\frac{1-p^{-r}}{p-1}p^{2s}.$
All higher levels are $B$-dominated, so $v_{n}(r)=p^{n-2s}v_{2s}(r)$. ∎
We can now prove Conjecture 27(a).
Theorem 8.11 (Conjecture 27(a)).
Let $k\geq 1$ satisfy $p\mid k$. Then
$\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\frac{1-p^{-r}}{p-1}k+O(1).$
More precisely, if $k=\sum_{i\geq 0}k_{i}p^{i}$, then
$\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\sum_{i\geq 0}k_{i}v_{i}(r),$
and the error with respect to the linear function $\frac{1-p^{-r}}{p-1}k$ is uniformly bounded.
Proof.
If $k$ is a power of $p$, this is Corollary 8.10. Otherwise $p\mid k$ and Statement 8.4 gives
$\operatorname{ord}_{p}(a_{k}(r))=\Lambda_{r}(k)=\sum_{i\geq 0}k_{i}v_{i}(r).$
For all sufficiently large $i$, namely $i\geq 2s$, Corollary 8.10 gives
$v_{i}(r)=\frac{1-p^{-r}}{p-1}p^{i}.$
Therefore the difference
$\Lambda_{r}(k)-\frac{1-p^{-r}}{p-1}k$
comes only from the finitely many digits below level $2s$, hence is bounded independently of $k$. ∎
The radius statement Conjecture 27(b) is already available from the general theorem of Fu and Nie [4], so the higher-fiber analysis in the present paper is complete at this point.